### v3

# Position Paper: An Integrated Perspective on Data, Metrics, and Methodology for Deep Time-Series Forecasting

Abstract
Deep time-series forecasting plays an integral role in numerous practical applications. However, existing research fall short by focusing narrowly on either neural architecture designs for long-term point forecasts or probabilistic models for short-term scenarios. By proposing a comprehensive framework, facilitated by a novel tool, ProbTS, that integrates diverse data scenarios, evaluation metrics, and methodological focuses, we aim to transcend the limitations of current forecasting practices. Rigorous experimentation uncovers pivotal insights, including the supreme importance of aligning forecasting methodologies with the unique characteristics of the data; the necessity of a broad spectrum of metrics for accurately assessing both point and distributional forecasts; and the challenges inherent in adapting existing forecasting methods to a wider range of scenarios. These findings not only challenge conventional approaches but also illuminate promising avenues for future research, suggesting a more nuanced and effective strategy for advancing the field of deep time-series forecasting. [^1] The ProbTS tool will be open-sourced.
time-series forecasting, point forecasting, probabilistic forecasting, neural networks
1 Introduction
Deep time-series forecasting serves a crucial role in a myriad of practical applications, from traffic flow forecasting (Lv et al., 2014) and renewable energy forecasting (Wang et al., 2019), to diverse forecasting demands in retail (Böse et al., 2017), finance (Hou et al., 2021), and climate (Mudelsee, 2019). Each of these domains presents unique data characteristics, such as trend movements (Taylor & Letham, 2018), periodic patterns (Smyl, 2020), complicated mixtures of global seasonality and local variations (Fan et al., 2022), as well as intricate distributional dependencies (Rasul et al., 2021b). These data characteristics collectively pose substantial challenges for accurate forecasting, and thereby necessitate the development of sophisticated modeling designs. In addition to accurate point forecasts, modern decision-making processes often require robust distributional forecasts to effectively manage uncertainty (Gneiting & Katzfuss, 2014; Hyndman & Athanasopoulos, 2018). These collective requirements emphasize the significance for methodologies that are not only versatile across various forecasting scenarios but also capable of delivering both point and distributional forecasts with high fidelity.
However, these practical needs also highlight a distinct gap in existing research for deep time-series forecasting, which tends to bifurcate into two streams: one focusing on neural architecture designs optimized for point forecasting in long-term scenarios, and the other on advancing probabilistic forecasting, primarily in short-term situations.
Particularly, one stream of research, exemplified by studies such as (Zhou et al., 2021; Wu et al., 2021; Liu et al., 2022; Zhang et al., 2022; Challu et al., 2023; Wu et al., 2023; Nie et al., 2023), primarily explores neural architecture designs tailored for time-series forecasting. These studies, however, mostly focused on point forecasts, and their architectural designs are largely driven by long-term forecasting scenarios that exhibit strong trending and seasonal patterns. Therefore, it remains unclear how these advancements in neural architecture designs can be effectively extended to distributional forecasts, and whether their specialized designs, which are closely tied to certain data patterns, maintain their effectiveness across varied data scenarios.
On the other hand, another line of research, including works like (Rasul et al., 2021b, a; Tashiro et al., 2021; Bilos et al., 2023; Kollovieh et al., 2023), emphasizes the adaption of deep generative modeling (Dinh et al., 2017; Papamakarios et al., 2017; Ho et al., 2020) to probabilistic time-series forecasting. Despite significant progress in characterizing complex data distributions without the need for a predefined closed-form prior distribution such as Gaussian (Salinas et al., 2019), these probabilistic models have primarily been developed and evaluated in short-term forecasting scenarios. Hence, it remains to be seen whether these designs can be effectively transitioned to long-term forecasting, and whether improvements in distributional forecasts also yield benefits in point forecasts, especially when compared to the tailored architectural designs mentioned earlier.
In a bid to systematically bridge this research gap, in this paper, we propose a holistic framework that integrates diverse data scenarios, comprehensive evaluation metrics, and methodological innovations. This approach enables a nuanced analysis of why certain forecasting methodologies outperform others in specific contexts and how they could be optimized to improve both point and distributional forecasting. To be more concrete, firstly, in terms of data, we encompass various data scenarios, spanning different domains and forecasting horizons. More significantly, we aim to quantify key data characteristics, elucidating why certain methodological choices converge or diverge. Secondly, concerning metrics, we adopt a comprehensive set of evaluation strategies, assessing the performance of both point and distributional forecasts. Lastly, we revisit a selection of representative studies from different research branches, affirming their advancements in specific methodological focuses and examining their performance in seldom evaluated scenarios. Through this integrated lens, we aspire to illuminate previously unaddressed questions, expose overlooked limitations of existing methods in certain situations, and reveal our novel findings that could potentially open up new avenues for future research.
To achieve these ends, we develop a tool, ProbTS, which incorporates customized data, evaluation, and model modules in alignment with the integrated perspective. Rigorous experimentation with this tool has led us to a number of intriguing insights, which are concisely presented below and explored in greater detail in Section 4.
Firstly, regarding the data, we have discovered that the distinctive characteristics of data in different forecasting scenarios primarily dictate the selection of methodology. We have identified several fundamental indicators that assist in distinguishing the distinct needs of diverse forecasting scenarios, such as data from various domains, short-term versus long-term forecasting, and point versus distributional forecasting. Notably, these indicators also facilitate our understanding of why certain methodological choices align or diverge in particular cases.
Secondly, concerning metrics, we have discerned that the use of a comprehensive suite of metrics is critical for both point and distributional forecasts. This conclusion stems from our observation that optimal performance in distributional metrics does not necessarily translate to optimal performance in point forecasts, and vice versa. Importantly, this suggests that focusing solely on optimizing one type of evaluation objective could result in sub-optimal performance according to other evaluation criteria.
Thirdly, in relation to the first research branch primarily dedicated to developing new architectures tailored for long-term point forecasting scenarios, we have recognized significant challenges and unresolved questions when adapting them to new scenarios, particularly those with distributional or short-term requirements. The integration of these architectures with advanced distribution estimation capabilities remains an open issue. Furthermore, the advantages of specific designs created for long-term scenarios diminish significantly in short-term cases. Another interesting observation is that all these methods consistently favor non-autoregressive decoding schemes over their autoregressive counterparts for multi-step forecasting, likely due to the pronounced trends in many long-term forecasting scenarios.
In terms of the second research branch, which focuses on developing advanced probabilistic forecasting capabilities, primarily in short-term scenarios, we have observed that extending existing probabilistic models to long-term scenarios presents significant challenges. Specifically, they often underperform those point forecasting methods even in producing distributional forecasts. Different from the first branch, these probabilistic methods exhibit a balanced preference for autoregressive and non-autoregressive decoding schemes. Our analyses reveal that autoregressive probabilistic methods tend to propagate more errors as the forecasting horizon increases, and non-autoregressive probabilistic methods demand substantial computational resources, both in terms of memory and time, as the forecasting horizon expands significantly. Interestingly, we also discovered that autoregressive probabilistic methods exhibit significant advantages in dealing with strong seasonality, even in long-term forecasting scenarios. This suggests that autoregressive decoding holds substantial potential for long-term forecasting, provided we can effectively mitigate the error-propagation effect amplified by pronounced trending.
Lastly, by leveraging our versatile tool that facilitates a deeper understanding of the interplay between data characteristics, forecasting performance, and methodological choices, we underscore several under-explored yet valuable directions that not only illuminate the path forward for academic research but also hold profound implications for practical applications in a multitude of sectors.
2 Related Work
As outlined in the introduction, deep time-series forecasting can be divided into two primary branches. This section provides a more thorough review of these branches, as well as a brief overview of existing forecasting toolkits. Besides, we include a refined comparison in Appendix A.1.
Neural Architecture Designs in Time-series Forecasting.
Significant research efforts have been devoted to improving neural architecture designs for time-series forecasting, with contributions ranging from extensions of multi-layer perceptrons (Oreshkin et al., 2020; Zhang et al., 2022) to customized recurrent or convolutional neural networks (Lai et al., 2018; LIU et al., 2022), and Transformer-based variants (Vaswani et al., 2017; Zhou et al., 2022). These studies typically employ non-autoregressive decoding schemes and focus on long-term forecasting scenarios, often characterized by strong trend and seasonality patterns. However, they predominantly limit themselves to point forecasts, capturing only the average future variations. While approaches like quantile regression (Wen et al., 2017; Lim et al., 2021) can mitigate this limitation, they still cannot replace the inherent capture of the data distribution.
Probabilistic Estimation in Time-series Forecasting.
In contrast, the second branch of research, known as deep probabilistic time-series forecasting, focuses on leveraging deep neural networks to capture the complex data distribution of future time series. These studies span from early approaches using pre-defined likelihood functions (Rangapuram et al., 2018; Salinas et al., 2020), to Gaussian copulas (Salinas et al., 2019; Drouin et al., 2022), and to more recent methods exploiting advanced deep generative models (Rasul et al., 2021b; Bilos et al., 2023). While providing probabilistic forecasts, they mainly evaluate short-term forecasting scenarios. Interestingly, unlike the first branch, this branch utilizes both autoregressive (Rasul et al., 2021b, a) and non-autoregressive (Tashiro et al., 2021; Bilos et al., 2023; Kollovieh et al., 2023) decoding schemes. Furthermore, this branch tends to use general neural architectures instead of customizing them. While a few methods (Tashiro et al., 2021; Li et al., 2022; Bergsma et al., 2022) incorporate customized network designs, more studies (de Bézenac et al., 2020; Rasul et al., 2021b, a; Bilos et al., 2023; Drouin et al., 2022) leverage standardized neural networks to encode time-series representations.
Toolkits for Time-series Forecasting.
We observe a plethora of toolkits that have been developed for time-series forecasting. These range from those primarily designed for point forecasting, such as Prophet (Taylor & Letham, 2018), sktime (Löning et al., 2019), tsai (Oguiza, 2022), and TSlib (Wu et al., 2023), to others that incorporate probabilistic forecasting, including GluonTS (Alexandrov et al., 2020), PyTorchTS (Rasul et al., 2021b), PyTorchForecasting222github.com/jdb78/pytorch-forecasting, and NeuralForecast333github.com/Nixtla/neuralforecast. In the development of ProbTS, we drew upon these existing tools, particularly PyTorchTS and GluonTS. However, ProbTS distinguishes itself by its specialized design towards the integrated perspective proposed in this paper, encompassing modularized designs across data, methodology, and evaluation components. Furthermore, we incorporate more state-of-the-art forecasting models in ProbTS, by migrating their official implementations and conducting extensive hyper-parameter tuning to ensure optimal performance.
3 The ProbTS Tool
This section offers a concise overview of the ProbTS tool’s design and implementation. The core modules and the primary pipeline of ProbTS are depicted in Figure 1.
Figure 1: An overview of ProbTS.
Data.
We aggregate publicly accessible datasets used for both short-term and long-term forecasting. Initial data visualization analyses reveal that the data domains and forecasting horizons significantly influence specific data characteristics within a given forecasting horizon. For instance, many long-term forecasting scenarios exhibit clear trend and seasonality patterns within a forecasting window, while numerous short-term forecasting cases display irregular variations within a short sliding window. Consequently, we have developed quantified indicators, such as trend and seasonality strengths, along with non-Gaussianity to indicate the complexity of data distribution within a forecasting window. Detailed information about dataset statistics, visualization analyses, and quantified measures can be found in Appendix A.2A.3A.4, and A.5. The quantified measurements for all forecasting scenarios are compiled in Table 1.
Metrics.
ProbTS incorporates a broad range of evaluation metrics to enable a thorough assessment of both point and distributional forecasts. These metrics are elaborated in detail in Appendix A.6. In this paper, we primarily use the normalized mean absolute error (NMAE) for point forecasts and the continuous ranked probability score (CRPS) for distributional forecasts to succinctly communicate the critical insights discovered. It is noteworthy that some methods reproduced in ProbTS, their original papers reported certain point forecast metrics before de-normalizing forecasts to the initial scale (Zeng et al., 2023; Wu et al., 2023; Nie et al., 2023) or primarily reveal aggregated distributional metrics over all time-series variates, namely CRPS-sum (Salinas et al., 2019; Rasul et al., 2021a, b). We have diligently ensured our reproduced results align with their reported results and utilized the unified metrics in this study to offer a comprehensive and fair comparison.
Model.
The design of our model module is modularized to accommodate various methodological focuses, such as point versus distributional forecasts, customized versus general neural architectures for encoding time-series representations, and autoregressive versus non-autoregressive decoding schemes. This design enables us to easily integrate with existing methods by selecting different routes and configuring specific components appropriately. Specifically, our design is grounded on the following formulation for time-series forecasting that considers different forecasting paradigms and decoding schemes.
Formulation.
We denote an element of a multivariate time series as $x^{k}_{t}\in\mathbb{R}$, where $k$ represents the variate index and $t$ denotes the time index. At time step $t$, we have a multivariate vector $\bm{x}_{t}\in\mathbb{R}^{K}$. Each $x^{k}_{t}$ is associated with covariates $\bm{c}_{t}^{k}\in\mathbb{R}^{N}$, which encapsulates auxiliary information about the observations. Given a length-$T$ forecast horizon a length-$L$ observation history $\bm{x}_{t-L:t}$ and corresponding covariates $\bm{c}_{t-L:t}$, the objective in time series forecasting is to generate the vector of future values $\bm{x}_{t+1:t+T}$. In ProbTS, we decouple a model into an encoder $f_{\phi}$ and a forecaster $p_{\theta}$. An encoder is tasked with generating expressive hidden states $\bm{h}\in\mathbb{R}^{D}$. Under autoregressive decoding scheme, encoder forecasts variates using their past values, which can be formulated as $\bm{h}_{t}=f_{\phi}(\bm{x}_{t-1},\bm{c}_{t},\bm{h}_{t-1})$. Under the non-autoregressive scheme, the encoder generates all the forecasts in one step, which can be expressed as $\bm{h}_{t+1:t+T}=f_{\phi}(\bm{x}_{t-L:t},\bm{c}_{t-L:t+T})$. A forecaster $p_{\theta}$ is employed either to directly estimate point forecasts as $\hat{\bm{x}}_{t}=p_{\theta}(\bm{h}_{t})$, or to perform sampling based on the estimated probabilistic distributions as $\hat{\bm{x}}_{t}\sim p_{\theta}(\bm{x}_{t}|\bm{h}_{t})$.
Table 1: This table presents a quantitative assessment of the inherent characteristics for all forecasting scenarios, each corresponding to a dataset with a specific forecasting horizon. We use the suffixes ”-S” and ”-L” to differentiate between short-term and long-term scenarios. Quantified indicators encompass trend and seasonality strengths, as well as non-Gaussianity, where a higher value signifies a greater deviation from a Gaussian distribution.
Dataset-Horizon Exchange-S Solar-S Electricity-S Traffic-S Wikipedia-S ETTm1-L ETTm2-L
Trend 0.9982 0.1688 0.6443 0.2880 0.5253 0.9462 0.9770
Seasonality 0.1256 0.8592 0.8323 0.6656 0.2234 0.0105 0.0612
Non-Gaussianity 0.2967 0.5004 0.3579 0.2991 0.2751 0.0833 0.1701
Dataset-Horizon ETTh1-L ETTh2-L Electricity-L Traffic-L Weather-L Exchange-L ILI-L
Trend 0.7728 0.9412 0.6476 0.1632 0.9612 0.9978 0.5438
Seasonality $F_{S}$ 0.4772 0.3608 0.8344 0.6798 0.2657 0.1349 0.6075
Non-Gaussianity 0.0719 0.1422 0.1533 0.1378 0.1727 0.1082 0.1112
4 Analyses
Table 2: Results ($\textrm{mean}_{\textrm{std}}$) on short-term forecasting scenarios, each containing five independent runs with different seeds.
Model Exchange Rate Solar Electricity Traffic Wikipedia
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
Glob. mean $0.188$ $0.188$ $1.403$ $1.403$ $0.412$ $0.412$ $0.540$ $0.540$ $0.577$ $0.577$
Batch mean $0.012$ $0.012$ $1.244$ $1.244$ $0.365$ $0.365$ $0.503$ $0.503$ $0.336$ $0.336$
Linear $0.012_{.001}$ $0.012_{.001}$ $0.704_{.036}$ $0.704_{.036}$ $0.138_{.009}$ $0.138_{.009}$ $0.327_{.032}$ $0.327_{.032}$ $0.874_{.151}$ $0.874_{.151}$
GRU $0.013_{.002}$ $0.013_{.002}$ $0.594_{.144}$ $0.594_{.144}$ $0.134_{.009}$ $0.134_{.009}$ $0.193_{.002}$ $0.193_{.002}$ $0.394_{.013}$ $0.394_{.013}$
Transformer $0.016_{.001}$ $0.016_{.001}$ $0.538_{.066}$ $0.538_{.066}$ $0.115_{.005}$ $0.115_{.005}$ $0.204_{.006}$ $0.204_{.006}$ $0.408_{.011}$ $0.408_{.011}$
N-HiTS $0.012_{.000}$ $0.012_{.000}$ $0.572_{.020}$ $0.572_{.020}$ $0.074_{.003}$ $0.074_{.003}$ $0.193_{.002}$ $0.193_{.002}$ $0.332_{.011}$ $0.332_{.011}$
NLinear $0.010_{.000}$ $0.010_{.000}$ $0.560_{.002}$ $0.560_{.002}$ $0.083_{.002}$ $0.083_{.002}$ $0.233_{.001}$ $0.233_{.001}$ $0.321_{.001}$ $0.321_{.001}$
DLinear $0.012_{.001}$ $0.012_{.001}$ $0.547_{.009}$ $0.547_{.009}$ $0.076_{.003}$ $0.076_{.003}$ $0.250_{.002}$ $0.250_{.002}$ $0.412_{.001}$ $0.412_{.001}$
PatchTST $0.010_{.000}$ $0.010_{.000}$ $0.496_{.002}$ $0.496_{.002}$ $0.067_{.001}$ $0.067_{.001}$ $0.202_{.001}$ $0.202_{.001}$ $0.257_{.001}$ $0.257_{.001}$
TimesNet $0.011_{.001}$ $0.011_{.001}$ $0.507_{.019}$ $0.507_{.019}$ $0.071_{.002}$ $0.071_{.002}$ $0.205_{.002}$ $0.205_{.002}$ $0.304_{.002}$ $0.304_{.002}$
GRU NVP $0.016_{.003}$ $0.020_{.003}$ $0.396_{.021}$ $0.507_{.022}$ $0.055_{.002}$ $0.073_{.003}$ $0.161_{.006}$ $0.203_{.009}$ $0.282_{.003}$ $0.330_{.003}$
GRU MAF $0.015_{.001}$ $0.020_{.001}$ $0.386_{.026}$ $0.492_{.027}$ $0.051_{.001}$ $0.067_{.001}$ $0.131_{.006}$ $0.165_{.009}$ $0.281_{.004}$ $0.337_{.005}$
Trans MAF $0.011_{.001}$ $0.014_{.001}$ $0.400_{.022}$ $0.503_{.022}$ $0.054_{.004}$ $0.071_{.005}$ $0.129_{.004}$ $0.165_{.006}$ $0.289_{.008}$ $0.344_{.008}$
TimeGrad $0.011_{.001}$ $0.014_{.002}$ $0.359_{.011}$ $0.445_{.023}$ $0.052_{.001}$ $0.067_{.001}$ $0.164_{.091}$ $0.201_{.115}$ $0.272_{.008}$ $0.327_{.011}$
CSDI $0.008_{.000}$ $0.011_{.000}$ $0.366_{.005}$ $0.484_{.008}$ $0.050_{.001}$ $0.065_{.001}$ $0.146_{.012}$ $0.176_{.013}$ $0.219_{.006}$ $0.259_{.009}$
Table 3: Results ($\textrm{mean}_{\textrm{std}}$) on long-term forecasting scenarios, each containing five independent runs with different seeds. The input sequence length is set to 36 for the ILI dataset and 96 for the others. Due to the excessive time and memory consumption of CSDI in producing long-term forecasts, its results are unavailable in some datasets.
Model pred DLinear PatchTST GRU NVP TimeGrad CSDI
len CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
ETTm1 96 $0.282_{.002}$ $0.282_{.002}$ $0.272_{.001}$ $0.272_{.001}$ $0.383_{.053}$ $0.488_{.058}$ $0.522_{.105}$ $0.645_{.129}$ $0.236_{.006}$ $0.308_{.005}$
192 $0.309_{.004}$ $0.309_{.004}$ $0.295_{.001}$ $0.295_{.001}$ $0.396_{.030}$ $0.514_{.042}$ $0.603_{.092}$ $0.748_{.084}$ $0.291_{.025}$ $0.377_{.026}$
336 $0.338_{.008}$ $0.338_{.008}$ $0.323_{.001}$ $0.323_{.001}$ $0.486_{.032}$ $0.630_{.029}$ $0.601_{.028}$ $0.759_{.015}$ $0.322_{.033}$ $0.419_{.042}$
720 $0.387_{.006}$ $0.387_{.006}$ $0.353_{.001}$ $0.353_{.001}$ $0.546_{.036}$ $0.707_{.050}$ $0.621_{.037}$ $0.793_{.034}$ $0.448_{.038}$ $0.578_{.051}$
ETTm2 96 $0.138_{.000}$ $0.138_{.000}$ $0.132_{.001}$ $0.132_{.001}$ $0.319_{.044}$ $0.413_{.059}$ $0.427_{.042}$ $0.525_{.047}$ $0.115_{.009}$ $0.146_{.012}$
192 $0.163_{.003}$ $0.163_{.003}$ $0.157_{.001}$ $0.157_{.001}$ $0.326_{.025}$ $0.427_{.033}$ $0.424_{.061}$ $0.530_{.060}$ $0.147_{.008}$ $0.189_{.012}$
336 $0.188_{.001}$ $0.188_{.001}$ $0.176_{.000}$ $0.176_{.000}$ $0.449_{.145}$ $0.580_{.169}$ $0.469_{.049}$ $0.566_{.047}$ $0.190_{.018}$ $0.248_{.024}$
720 $0.219_{.003}$ $0.219_{.003}$ $0.205_{.001}$ $0.205_{.001}$ $0.561_{.273}$ $0.749_{.385}$ $0.470_{.054}$ $0.561_{.044}$ $0.239_{.035}$ $0.306_{.040}$
ETTh1 96 $0.352_{.011}$ $0.352_{.011}$ $0.328_{.003}$ $0.328_{.003}$ $0.379_{.030}$ $0.481_{.037}$ $0.455_{.046}$ $0.585_{.058}$ $0.437_{.018}$ $0.557_{.022}$
192 $0.393_{.001}$ $0.393_{.001}$ $0.359_{.002}$ $0.359_{.002}$ $0.425_{.019}$ $0.531_{.018}$ $0.516_{.038}$ $0.680_{.058}$ $0.496_{.051}$ $0.625_{.065}$
336 $0.419_{.007}$ $0.419_{.007}$ $0.384_{.002}$ $0.384_{.002}$ $0.458_{.054}$ $0.580_{.064}$ $0.512_{.026}$ $0.666_{.047}$ $0.454_{.025}$ $0.574_{.026}$
720 $0.502_{.029}$ $0.502_{.029}$ $0.397_{.002}$ $0.397_{.002}$ $0.502_{.039}$ $0.643_{.046}$ $0.523_{.027}$ $0.672_{.015}$ $0.528_{.012}$ $0.657_{.014}$
ETTh2 96 $0.211_{.027}$ $0.211_{.027}$ $0.177_{.000}$ $0.177_{.000}$ $0.432_{.141}$ $0.548_{.158}$ $0.358_{.026}$ $0.448_{.031}$ $0.164_{.013}$ $0.214_{.018}$
192 $0.238_{.028}$ $0.238_{.028}$ $0.201_{.001}$ $0.201_{.001}$ $0.625_{.170}$ $0.766_{.223}$ $0.457_{.081}$ $0.575_{.089}$ $0.226_{.018}$ $0.294_{.027}$
336 $0.284_{.008}$ $0.284_{.008}$ $0.240_{.001}$ $0.240_{.001}$ $0.793_{.319}$ $0.942_{.408}$ $0.481_{.078}$ $0.606_{.095}$ $0.274_{.022}$ $0.353_{.028}$
720 $0.307_{.000}$ $0.307_{.000}$ $0.252_{.000}$ $0.252_{.000}$ $0.539_{.090}$ $0.688_{.161}$ $0.445_{.016}$ $0.550_{.018}$ $0.302_{.040}$ $0.382_{.030}$
Electricity 96 $0.090_{.001}$ $0.090_{.001}$ $0.086_{.001}$ $0.086_{.001}$ $0.094_{.003}$ $0.118_{.003}$ $0.096_{.002}$ $0.119_{.003}$ $0.153_{.137}$ $0.203_{.189}$
192 $0.095_{.001}$ $0.095_{.001}$ $0.092_{.001}$ $0.092_{.001}$ $0.097_{.002}$ $0.121_{.003}$ $0.100_{.004}$ $0.124_{.005}$ $0.200_{.094}$ $0.264_{.129}$
336 $0.104_{.000}$ $0.104_{.000}$ $0.100_{.000}$ $0.100_{.000}$ $0.099_{.001}$ $0.123_{.001}$ $0.102_{.007}$ $0.126_{.008}$
720 $0.122_{.001}$ $0.122_{.001}$ $0.116_{.000}$ $0.116_{.000}$ $0.114_{.013}$ $0.144_{.017}$ $0.108_{.003}$ $0.134_{.004}$
Traffic 96 $0.356_{.009}$ $0.356_{.009}$ $0.248_{.001}$ $0.248_{.001}$ $0.187_{.002}$ $0.231_{.003}$ $0.202_{.004}$ $0.234_{.006}$
192 $0.346_{.009}$ $0.346_{.009}$ $0.245_{.001}$ $0.245_{.001}$ $0.192_{.001}$ $0.236_{.002}$ $0.208_{.003}$ $0.239_{.004}$
336 $0.350_{.008}$ $0.350_{.008}$ $0.257_{.002}$ $0.257_{.002}$ $0.201_{.004}$ $0.248_{.006}$ $0.213_{.003}$ $0.246_{.003}$
720 $0.365_{.009}$ $0.365_{.009}$ $0.266_{.001}$ $0.266_{.001}$ $0.211_{.004}$ $0.264_{.006}$ $0.220_{.002}$ $0.263_{.001}$
Weather 96 $0.112_{.001}$ $0.112_{.001}$ $0.087_{.002}$ $0.087_{.002}$ $0.116_{.013}$ $0.145_{.017}$ $0.130_{.017}$ $0.164_{.023}$ $0.068_{.008}$ $0.087_{.012}$
192 $0.122_{.001}$ $0.122_{.001}$ $0.090_{.001}$ $0.090_{.001}$ $0.122_{.021}$ $0.147_{.025}$ $0.127_{.019}$ $0.158_{.024}$ $0.068_{.006}$ $0.086_{.007}$
336 $0.130_{.002}$ $0.130_{.002}$ $0.092_{.002}$ $0.092_{.002}$ $0.128_{.011}$ $0.160_{.012}$ $0.130_{.006}$ $0.162_{.006}$ $0.083_{.002}$ $0.098_{.002}$
720 $0.144_{.001}$ $0.144_{.001}$ $0.094_{.001}$ $0.094_{.001}$ $0.110_{.004}$ $0.135_{.008}$ $0.113_{.011}$ $0.136_{.020}$ $0.087_{.003}$ $0.102_{.005}$
Exchange 96 $0.024_{.000}$ $0.024_{.000}$ $0.023_{.000}$ $0.023_{.000}$ $0.071_{.006}$ $0.091_{.009}$ $0.068_{.003}$ $0.079_{.002}$ $0.028_{.003}$ $0.036_{.005}$
192 $0.035_{.000}$ $0.035_{.000}$ $0.034_{.000}$ $0.034_{.000}$ $0.068_{.004}$ $0.087_{.005}$ $0.087_{.013}$ $0.100_{.019}$ $0.045_{.003}$ $0.058_{.005}$
336 $0.048_{.001}$ $0.048_{.001}$ $0.048_{.000}$ $0.048_{.000}$ $0.072_{.002}$ $0.091_{.002}$ $0.074_{.009}$ $0.086_{.008}$ $0.060_{.004}$ $0.076_{.006}$
720 $0.075_{.002}$ $0.075_{.002}$ $0.072_{.000}$ $0.072_{.000}$ $0.079_{.009}$ $0.103_{.009}$ $0.099_{.015}$ $0.113_{.016}$ $0.143_{.020}$ $0.173_{.020}$
ILI 24 $0.213_{.038}$ $0.213_{.038}$ $0.169_{.005}$ $0.169_{.005}$ $0.257_{.003}$ $0.283_{.001}$ $0.275_{.047}$ $0.296_{.044}$ $0.250_{.013}$ $0.263_{.012}$
36 $0.230_{.015}$ $0.230_{.015}$ $0.156_{.005}$ $0.156_{.005}$ $0.281_{.004}$ $0.307_{.007}$ $0.272_{.057}$ $0.298_{.048}$ $0.285_{.010}$ $0.298_{.011}$
48 $0.221_{.009}$ $0.221_{.009}$ $0.156_{.008}$ $0.156_{.008}$ $0.288_{.008}$ $0.314_{.009}$ $0.295_{.033}$ $0.320_{.025}$ $0.285_{.036}$ $0.301_{.034}$
60 $0.230_{.013}$ $0.230_{.013}$ $0.147_{.003}$ $0.147_{.003}$ $0.307_{.005}$ $0.333_{.005}$ $0.295_{.083}$ $0.325_{.068}$ $0.283_{.012}$ $0.299_{.013}$
[Long-term forecasting (NMAE).] Refer to caption [Long-term forecasting (CRPS).] Refer to caption     Refer to caption
[Short-term forecasting (NMAE).] Refer to caption [Short-term forecasting (CRPS).] Refer to caption     Refer to caption
Figure 2: We present plots of the metrics against non-Gaussianity for various models across all forecasting scenarios. In each subplot, a vertical line of points represents one forecasting scenario, denoted by a sequence number at the top. We distinguish between point and probabilistic forecasting methods using different colors, and we use unique markers to indicate various models.
[CRPS w.r.t. forecasting horizon.] Refer to caption     [CRPS w.r.t. trend strength.] Refer to caption
[CRPS w.r.t. seasonality strength.] Refer to caption     [The CRPS Gap, computed as TimeGrad.CRPS - CSDI.CRPS, w.r.t. trend and seasonal strengths.] Refer to caption
Figure 3: We analyze CRPS in relation to various factors to scrutinize the strengths and weaknesses of autoregressive probabilistic methods. Subplot (a), based on ETTm and ILI datasets, illustrates the error amplification that occurs with increased prediction length. Subplots (b) and (c) encompass all long-term forecasting scenarios, aiming to explore the relationship between CRPS and the strengths of trends and seasonality, respectively. Subplot (d) includes all forecasting scenarios, with a lighter and smaller circle denoting a preference for autoregressive methods over non-autoregressive ones.
Utilizing ProbTS, we benchmarked selected methods in both short-term and long-term forecasting scenarios. We evaluated these methods using both point and distributional forecast metrics. To ensure the reliability of our results and minimize randomness, we performed each experiment five times with different seeds for parameter initialization. The results for short-term and long-term forecasting are presented in Table 2 and Table 3, respectively.
Selected Methods for Comparison.
We have chosen representative methods from various research branches for comparison. Our selection is based on a comprehensive consideration of factors such as state-of-the-art performance, reproducibility, as well as the simplicity, universality, and novelty of the method designs. Adhering to these principles, we have selected PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023), N-HiTS (Challu et al., 2023), LTSF-Linear (NLinear and DLinear) (Zeng et al., 2023) from the research line that concentrates on developing customized neural architectures for long-term point forecasting. In addition, we have chosen advanced probabilistic forecasting methods for comparison, including those built upon normalizing flows such as GRU NVP, GRU MAF, and Trans MAF (Rasul et al., 2021b), and diffusion models such as TimeGrad (Rasul et al., 2021a) and CSDI (Tashiro et al., 2021). We also include general network architectures, such as Linear, GRU (Chung et al., 2014), and Transformer (Vaswani et al., 2017), to demonstrate the additional benefits brought by advanced architecture designs or probabilistic estimations. Simple non-parametric baselines, such as calculating the global mean or batch mean as all forecasts, are also included to assess the predictability of a forecasting scenario.
4.1 Discrepancies in Point and Distributional Metrics.
Our analysis revealed that optimal performance in distributional metrics does not guarantee optimal performance in point forecasts. This pattern emerged in several instances, notably in Table 3. For example, CSDI excels in the CRPS metric but falls short in NMAE across multiple datasets, such as ETTm1, ETTm2, and ETTh2. These observations underscore the necessity of a comprehensive evaluation utilizing both NMAE and CRPS metrics to gain a holistic understanding of performance in both point and distributional forecasting. It is worth highlighting that these two metrics become equivalent for point forecasting methods when a point forecast is converted into a degenerate distribution, centered on the forecasted point, for CRPS evaluation (Gneiting & Raftery, 2007).
4.2 Examination of Customized Neural Architectures
Our findings affirm the progress in this area of research. For instance, as observed in Table 3, the state-of-the-art architecture, PatchTST, consistently surpasses other models, especially when assessed using the NMAE metric. The only exception is on the Traffic dataset, where PatchTST is outperformed by GRU NVP and TimeGrad. This discrepancy is primarily attributable to the different decoding schemes utilized, which will be elaborated in Section 4.3.
Diminishing Advantages of Customized Architectures in Short-term Forecasting Scenarios.
Interestingly, the significant superiority of PatchTST diminishes in short-term forecasting scenarios. This trend is evident when comparing Table 3 and 2. More intuitively, Figures 22 display the relative strengths of point and probabilistic forecasting methods in terms of NMAE, and correlates their NMAE values with our quantified indicator, non-Gaussianity, used for assessing data distribution complexity. We further observe that as non-Gaussianity increases, signifying a more complex data distribution, a noticeable performance gap emerges. This suggests that the reducing superiority of these customized architectures may be due to their limited capability to handle complex data distributions.
Challenges in Integrating Advanced Distribution Estimation Capabilities with These Architectures.
Regrettably, the integration of these architectures with advanced distribution estimation capabilities remains a challenging task. As will be discussed subsequently, non-autoregressive probabilistic methods encounter significant efficiency issues in long-term forecasting. Intriguingly, all these architectures select a non-autoregressive decoding scheme. We will delve into the advantages and disadvantages of autoregressive and non-autoregressive schemes in the next section.
4.3 Assessment of Probabilistic Forecasting Methods
At the outset, we validate the efficacy of these probabilistic methods in the short-term forecasting scenarios they were designed for. Evidence from Table 2 demonstrates that innovative methods such as CSDI and TimeGrad yield superior results in most cases.
However, in Long-term Forecasting Scenarios, Probabilistic Methods Fall Short, Even When Evaluated Using Distributional Metrics.
As depicted in Table 3, in a multitude of long-term forecasting scenarios, customized neural architectures, created specifically with point forecasting objectives in mind, decidedly outperform probabilistic methods, even when evaluated using distributional metrics such as CRPS. This suggests that existing probabilistic methods may encounter significant challenges when applied to long-term forecasting, leaving substantial room for improvement in this domain. To further illustrate this, we compare the CRPS performance of these models with the non-Gaussianity in both long-term and short-term scenarios, as depicted in Figure 2 and 2, respectively. While non-Gaussianity does provide some insights—its values are relatively lower in long-term scenarios compared to short-term contexts—these explanations are not comprehensive. This is because it remains unclear to what extent a customized neural model, coupled with effective probabilistic estimations, can perform. To delve deeper, we further examine other unique aspects of the existing probabilistic forecasting methods in search of potential clues.
The Dichotomy of Decoding Scheme Preferences.
Upon further investigation, we find that these probabilistic forecasting methods exhibit a balanced preference for both autoregressive and non-autoregressive decoding schemes. For instance, TimeGrad employs an autoregressive decoding scheme, whereas CSDI utilizes a non-autoregressive decoding approach. This contrasts starkly with the aforementioned customized architectures, which solely opt for non-autoregressive decoding. These two types of decoding schemes, however, confront distinctive challenges when applied to long-term probabilistic forecasting.
  • Autoregressive probabilistic methods, for example, TimeGrad, are prone to propagate more errors as the forecasting horizon increases or the trend strength enlarges. This phenomenon is evident in Figure 3 and 3, where we observe that the performance gap escalates with either extended prediction lengths or intensified trend strengths.
  • In contrast, non-autoregressive probabilistic methods, like CSDI, struggle with significant efficiency challenges, both in memory usage and time consumption, in long-term forecasting scenarios. This issue is further detailed in Appendix Section B.6. Furthermore, as indicated in Table 3, even for relatively small-scale datasets, such as ETTm2 and ETTh1, CSDI’s performance in long-term scenarios is less than optimal. This suggests that its learning efficacy is also compromised by the extension of the forecasting horizon.
The Unexpected Superiority of Autoregressive Decoding Scheme in Addressing Strong Seasonality.
Interestingly, the autoregressive decoding scheme exhibits a remarkable proficiency in dealing with strong seasonality. A prime example is the results for the Traffic dataset shown in Table 3, where both GRU NVP and TimeGrad outperform PatchTST, despite utilizing relatively standard neural architectures without any customization. In Figure 3, we further investigate this phenomenon by relating the strength of seasonality to the variance in performance gaps. Additionally, in Figure 3, we conduct a collective analysis of the interactions among CRPS Gap, trend, and seasonality strengths, where the CRPS Gap is defined as the difference between the CRPS scores of TimeGrad (autoregressive) and CSDI (non-autoregressive). A clear pattern emerges: as the strength of seasonality heightens, autoregressive methods gain more advantages. This observation suggests that despite its challenges, the autoregressive decoding scheme should not be dismissed in long-term forecasting. Its unique affinity with seasonality, possibly due to superior parameter efficiency, may regain significance if we can develop effective strategies to mitigate the issue of error propagation.
4.4 Future Directions
Drawing on our analyses, we propose several promising directions for future research. Firstly, the development of hybrid learning objectives that optimize both point and distributional forecasts represents a significant advancement. This strategy could foster models capable of delivering more comprehensive predictions, catering to a diverse range of practical applications that necessitate both accurate predictions and well-quantified uncertainties. Secondly, there is an urgent need for the effective integration of neural architecture design with advanced probabilistic estimation paradigms. Addressing this demand could lay the groundwork for more efficient and precise forecasting models. Particularly, overcoming the efficiency challenges associated with non-autoregressive decoding schemes and reducing error propagation in autoregressive probabilistic models, specifically in the presence of marked trends, could substantially enhance the models’ performance and practicality. Lastly, the exploration of the necessity and potential designs of customized architectures tailored for short-term forecasting could initiate a novel research field. Such exploration necessitates consideration of data characteristics beyond trends and seasonality.
4.5 Limitations
Despite the thoroughness of our study, there are several limitations. First, due to the vast number of methods available in this field, we could not include all of them in our analysis. Our selection is based on a set of criteria, but inevitably, some methods were left out. Second, while we have tried to incorporate a diverse range of datasets, there are certainly more datasets beyond those included in our study. To mitigate this limitation, we have conducted additional experiments on univariate time-series datasets, which are included in Appendix B.3. Finally, our study is primarily empirical in nature. While we have made significant efforts to conduct rigorous statistical analyses, we have not provided theoretical proofs for our results.
5 Conclusion
This paper presents an integrated perspective on data, metrics, and methodology in the field of deep time-series forecasting, facilitated by the ProbTS tool. Our findings not only provide valuable insights into the strengths and limitations of existing methods but also highlight several underexplored areas that merit future research. In addition, we expect that our insights could spur more explorations on these fields. We also hope the ProbTS tool can become a useful resource for the community, expediting future advancements in time-series forecasting.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A More Details on ProbTS
A.1 Additional Related Work
Table 4 presents a comparative summary of our approach, which adopts an integrated perspective, and representative studies from the existing literature.
Table 4: We provide a concise comparison between the methodologies presented in this paper and those from two distinct research branches. The comparison is based on data scenarios (short-term versus long-term forecasting), primary evaluation metrics (point versus distributional forecasts), and key methodological choices (general or customized neural architecture designs, and autoregressive or non-autoregressive decoding schemes).
Method Fore. Horizon Metrics Arch. Design Dec. Scheme
Short Long Point Distr. General Customized Autoreg. Non-autoreg.
N-BEATS (Oreshkin et al., 2020)
Autoformer (Wu et al., 2021)
Informer (Zhou et al., 2021)
Pyraformer (Liu et al., 2022)
N-HiTS (Challu et al., 2023)
LTSF-Linear (Zeng et al., 2023)
PatchTST (Nie et al., 2023)
TimesNet (Wu et al., 2023)
DeepAR (Salinas et al., 2020)
GP-copula (Salinas et al., 2019)
LSTM NVP (Rasul et al., 2021b)
LSTM MAF (Rasul et al., 2021b)
Trans MAF (Rasul et al., 2021b)
TimeGrad (Rasul et al., 2021a)
CSDI (Tashiro et al., 2021)
SPD (Bilos et al., 2023)
TSDiff (Kollovieh et al., 2023)
This Paper
A.2 Time-series Forecasting Datasets
Table 5 provides a summary of the public datasets employed in our study. These datasets have been sourced from recent research studies in the field of deep time-series forecasting.
Table 5: Dataset Summary.
Horizon Dataset #var. range freq. timesteps Description
Long-term ETTh1/h2 7 $\mathbb{R}^{+}$ H 17,420 Electricity transformer temperature per hour
ETTm1/m2 7 $\mathbb{R}^{+}$ 15min 69,680 Electricity transformer temperature every 15 min
Electricity 321 $\mathbb{R}^{+}$ H 26,304 Electricity consumption (Kwh)
Traffic 862 (0,1) H 17,544 Road occupancy rates
Exchange 8 $\mathbb{R}^{+}$ Busi. Day 7,588 Daily exchange rates of 8 countries
ILI 7 (0,1) W 966 Ratio of patients seen with influenza-like illness
Weather 21 $\mathbb{R}^{+}$ 10min 52,696 Local climatological data
Short-term Exchange 8 $\mathbb{R}^{+}$ Busi. Day 6,071 Daily exchange rates of 8 countries
Solar 137 $\mathbb{R}^{+}$ H 7,009 Solar power production records
Electricity 370 $\mathbb{R}^{+}$ H 5,833 Electricity consumption
Traffic 963 (0,1) H 4,001 Road occupancy rates
Wikipedia 2,000 $\mathbb{N}$ D 792 Page views of 2000 Wikipedia pages
A.3 Data Visualization
To provide a more tangible understanding of the different forecasting scenarios, we visualize time-series segments from both short-term and long-term forecasting datasets. The segments’ window size is determined by the specific forecasting setup.
In Figure 4, we present samples extracted from short-term forecasting scenarios. At this scale, the series primarily exhibit local variations, and the compact window size often obscures pronounced seasonal or trending patterns. However, these short-term scenarios may reveal irregularly varied patterns, suggesting a more complex underlying data distribution.
On the contrary, Figure 5 illustrates long-term forecasting scenarios. With extended forecasting horizons, as showcased in datasets like Traffic, Electricity, and ETT, the series display more pronounced seasonality and trends. These characteristics render the series more regular patterns in the long-term scenarios.
It’s important to note that these visualizations are not selectively chosen or ”cherry-picked”. We have depicted multiple time-series segments from various time steps, and the observed patterns remain consistent across these different instances.
Figure 4: We have sampled and visualized multiple time-series segments from the short-term forecasting datasets. The size of the segment window is set equal to the prediction horizon.
Figure 5: We have also sampled and visualized multiple time-series segments from the long-term forecasting datasets, where the size of the segment window matches the prediction horizon.
A.4 Quantifying Trend and Seasonality Strengths
Based on the intuition obtained from data visualization, we would like to quantify the strengths of trend and seasonality for a time-series segment with a predefined window size (corresponding to the prediction horizon length). Then we can quantify the trend and seasonality strengths at the dataset level by averaging over all time-series segments of a dataset.
To quantify the strengths of trend and seasonality for a fixed-length time-series segment, we draw upon methodologies outlined in the work of (Wang et al., 2006). In particular, we employed a time series decomposition model expressed as:
$y_{t}=T_{t}+S_{t}+R_{t},$
where $T_{t}$ represents the smoothed trend component, $S_{t}$ signifies the seasonal component, and $R_{t}$ denotes the remainder component. In order to obtain each component, we followed the STL decomposition approach  [^4] https://otexts.com/fpp2/stl.html.
In the case of strongly trended data, the variation within the seasonally adjusted data should considerably exceed that of the remainder component. Consequently, the ratio $\textrm{Var}(R_{t})/\textrm{Var}(T_{t}+R_{t})$ is expected to be relatively small. As such, the measure of trend strength can be formulated as:
$F_{T}=\max\left(0,1-\frac{\textrm{Var}(R_{t})}{\textrm{Var}(T_{t}+R_{t})}\right).$
The quantified trend strength, ranging from 00 to $1$, characterizes the degree of trend presence. Similarly, the evaluation of seasonal intensity employs the detrended data:
$F_{S}=\max\left(0,1-\frac{\textrm{Var}(R_{t})}{\textrm{Var}(S_{t}+R_{t})}\right).$
A series with $F_{S}$ near 00 indicates minimal seasonality, while strong seasonality is indicated by $F_{S}$ approaching $1$ due to the considerably smaller variance of $\textrm{Var}(R_{t})$ in comparison to $\textrm{Var}(S_{t}+R_{t})$.
Tables 1 depict the results for each dataset. Notably, the ETT datasets and the Exchange dataset manifest conspicuous trends, whereas the Electricity, Solar, and Traffic datasets showcase marked seasonality. Additionally, the Exchange dataset stands out with distinctive features. Figure 5 also illustrates that with shorter prediction windows, the Exchange dataset sustains comparatively minor fluctuations, almost forming a linear trajectory. This enables effective forecasting through a straightforward batch mean approach. As the forecasting horizon extends, the dataset appears a more pronounced trend while retaining minimal seasonality.
A.5 Quantifying Data Distribution Complexity
To differentiate between methods optimized for point or distributional forecasts, we aim to quantify the complexity of data distribution within a time-series segment whose window size equals the prediction horizon length. We propose that assessing non-Gaussianity, i.e., how closely the distribution of time-series values within that window resembles a Gaussian distribution, could serve as a meaningful measure. This is because point forecasting methods, optimized with mean squared loss, are essentially equivalent to probabilistic counterparts that include a Gaussian output head and employ maximum a posteriori estimation. This suggests that point forecasting methods inherently assume that time-series values adhere to a Gaussian distribution. In contrast, advanced probabilistic methods, which do not make prior assumptions about data distribution, can adapt to complex data distributions in a data-driven manner.
Hence, we use the Jensen–Shannon divergence (Nielsen, 2019) to measure the similarity between the actual value distribution of a time-series segment and a Gaussian distribution fitted to the observed values. Short-term datasets used a window size of 30, while long-term datasets used a size of 336. By averaging the calculated divergence values across all time-series segments of a dataset, we obtain a dataset-level measure of non-Gaussianity. A larger divergence value indicates a larger deviation from a Gaussian distribution in the data.
A.6 Evaluation Metrics
In ProbTS, we integrate an extensive variety of metrics that take into account both point and distributional forecasts, thereby providing a comprehensive and multifaceted assessment of forecasting models.
A.6.1 Metrics for Point Forecasts
Mean Absolute Error (MAE)
The Mean Absolute Error (MAE) quantifies the average absolute deviation between the forecasts and the true values. Since it averages the absolute errors, MAE is robust to outliers. Its mathematical formula is given by:
$\textrm{MAE}=\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{T}_{t=1}|x_{i,t}-\hat{x}_{i,t}|,$
where $K$ is the number of variates, $L$ is the length of series, $x_{i,t}$ and $\hat{x}_{i,t}$ denotes the ground-truth value and the predicted value, respectively. For multivariate time series, we also provide the aggregated version:
$\textrm{MAE}_{\textrm{sum}}=\frac{1}{T}\sum^{T}_{t=1}|x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t}|,$
where $x^{\textrm{sum}}_{t}$ and $\hat{x}^{\textrm{sum}}_{t}$ are the summation across the dimension $K$ of $x_{i,t}$ and $\hat{x}_{i,t}$, respectively.
Normalized Mean Absolute Error (NMAE)
The Normalized Mean Absolute Error (NMAE) is a normalized version of the MAE, which is dimensionless and facilitates the comparability of the error magnitude across different datasets or scales. The mathematical representation of NMAE is given by:
$\textrm{NMAE}=\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{T}_{t=1}\frac{|x_{i,t}-\hat{x}_{i,t}|}{|x_{i,t}|}.$
Its aggregated version is:
$\textrm{NMAE}_{\textrm{sum}}=\frac{1}{T}\sum^{T}_{t=1}\frac{|x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t}|}{|x^{\textrm{sum}}_{t}|}.$
Mean Squared Error (MSE)
The Mean Squared Error (MSE) is a quantitative metric used to measure the average squared difference between the observed actual value and forecasts. It is defined mathematically as follows:
$\textrm{MSE}=\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{L}_{t=1}(x_{i,t}-\hat{x}_{i,t})^{2}.$
For multivariate time series, we also provide the aggregated version:
$\textrm{MSE}_{\textrm{sum}}=\frac{1}{T}\sum^{L}_{t=1}(x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t})^{2}.$
Normalized Root Mean Squared Error (NRMSE)
The Normalized Root Mean Squared Error (NRMSE) is a normalized version of the Root Mean Squared Error (RMSE), which quantifies the average squared magnitude of the error between forecasts and observations, normalized by the expectation of the observed values. It can be formally written as:
$\textrm{NRMSE}=\frac{\sqrt{\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{L}_{t=1}(x_{i,t}-\hat{x}_{i,t})^{2}}}{\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{T}_{t=1}|x_{i,t}|}.$
For multivariate time series, we also provide the aggregated version:
$\textrm{NRMSE}_{\textrm{sum}}=\frac{\sqrt{\frac{1}{T}\sum^{L}_{t=1}(x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t})^{2}}}{\frac{1}{T}\sum^{T}_{t=1}|x^{\textrm{sum}}_{t}|}.$
A.6.2 Metrics for Distributional Forecasts
Continuous Ranked Probability Score (CRPS)
The Continuous Ranked Probability Score (CRPS) (Matheson & Winkler, 1976) quantifies the agreement between a cumulative distribution function (CDF) $F$ and an observation $x$, represented as:
[…]
where $\mathds{I}{x\leq z}$ denotes the indicator function, equating to one if $x\leq z$ and zero otherwise.
Being a proper scoring function, CRPS reaches its minimum when the predictive distribution $F$ coincides with the data distribution. When using the empirical CDF of $F$, denoted as $\hat{F}(z)=\frac{1}{n}\sum^{n}_{i=1}\mathds{I}\{X_{i}\leq z\}$, where $n$ represents the number of samples $X_{i}\sim F$, CRPS can be precisely calculated from the simulated samples of the conditional distribution $p_{\theta}(\bm{x}_{t}|\bm{h}_{t})$. In our practice, 100 samples are employed to estimate the empirical CDF.
For multivariate time series, the aggregate CRPS, denoted as $\textrm{CRPS}_{\textrm{sum}}$, is derived by summing across the $K$ time series, both for the ground-truth data and sampled data, and subsequently averaging over the forecasting horizon. Formally, it is represented as:
$\textrm{CRPS}_{\textrm{sum}}=\mathds{E}_{t}\left[\textrm{CRPS}\left(\hat{F}_{\textrm{sum}}(t),\sum^{K}_{i=1}x^{0}_{i,l}\right)\right].$
A.7 Implementation Details
ProbTSwas developed using PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019). During training, we sampled 100 batches per epoch and limited training to 50 epochs, using the CRPS metric for checkpointing. All experiments employed the Adam optimizer and were run on single NVIDIA Tesla V100 GPUs with CUDA 11.3. To enable evaluation of distribution-level metrics, we conducted 100 samplings to calculate metrics on the test set.
Following the most commonly adopted settings (Zeng et al., 2023; Nie et al., 2023; Wu et al., 2023), in the long-term forecasting context, all of the models are following the same experimental setup with prediction length $T\in\{24,36,48,60\}$ for ILI dataset and $T\in\{96,192,336,720\}$ for other datasets. Note that the lookback window here is 96 for all the models, to ensure a fair comparison. In the short-term forecasting context, the length of the lookback window is the same as the forecasting horizons, which are 30 for Exchange-S dataset and Wikipedia-S dataset, and 24 for the rest, the same as (Salinas et al., 2019).
Hyper-parameter Tuning
For a fair comparison, we conducted a comprehensive grid search for critical hyperparameters across all models in this study. Table 6 details the shared hyperparameters tuned within the ProbTSpipeline, along with those kept constant. Due to the vast array of model-specific hyperparameters, we present an example configuration in Table 7. Complete hyperparameter configurations for each model, identified through this process, will be made available in a public GitHub repository for transparency and reproducibility.
Table 6: Hyper-parameters values fixed or range searched in hyper-parameter tuning.
Hyper-parameter Value or Range Searched
learning rate [1e-4, 1e-3, 1e-2]
dropout [0, 0.1, 0.2]
batch_size [8, 16, 32, 64]
use_lags [True, False]
use_feat_idx_emb [True, False]
use_time_feat [True, False]
autoregressive [True, False]
scaler [Standard, Scaling, None]
limit_train_batches 100
num_samples 100
quantiles_num 20
Table 7: Hyperparameter settings for Electricity-S dataset.
Model Hyperparameter
DLinear learning_rate=0.01, kernel_size=3, f_hidden_size=40
PatchTST learning_rate=0.0001, stride=3, patch_len=6, n_layers=3, n_heads=8, dropout=0.1, kernel_size=3, f_hidden_size=32
TimesNet learning_rate=0.001, n_layers=2, num_kernels=6, top_k=5, f_hidden_size=64, d_ff=64
GRU NVP learning_rate=0.001, f_hidden_size=40, num_layers=2, n_blocks=3, hidden_size=100, conditional_length=200
GRU MAF learning_rate=0.001, f_hidden_size=40, num_layers=2, n_blocks=4, hidden_size=100, conditional_length=200
Trans MAF learning_rate=0.001, f_hidden_size=32, num_heads=8, n_blocks=4, hidden_size=100, conditional_length=200
TimeGrad learning_rate=0.001, f_hidden_size=128, num_layers=4, conditional_length=100, beta_end=0.1, diff_steps=100
CSDI learning_rate=0.001, channels=64, emb_time_dim=128, emb_feature_dim=16, num_steps=50, num_heads=8, n_layers=4
Appendix B Additional Results and Experiments
B.1 Impact of Data Scale
To further explore critical characteristics of time-series forecasting, we have examined the correlation between model performance gains, relative to the baseline model (GRU), and dataset dimensions, length, and volume (see Table 8). However, our analysis does not identify a significant correlation between these factors and model performance.
Table 8: The correlation coefficient between the data volume and the relative performance improvement compared to the baseline model (GRU).
Model DLinear PatchTST GRU NVP TimeGrad CSDI
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
# Var. 0.2422 0.2422 -0.2676 -0.2676 -0.1856 -0.2136 -0.1665 -0.1793 -0.2315 -0.2592
# Total timestep -0.1422 -0.1422 0.3821 0.3821 0.3072 0.3329 0.2860 0.2971 0.3542 0.3826
# Var. $\times$ Timestep 0.0162 0.0162 0.0166 0.0166 -0.0068 -0.0011 0.0082 0.0117 -0.0053 -0.0133
B.2 Statistical and Gradient Boosting Decision Tree Baselines
To enhance the empirical robustness of our study, we integrate classical statistical models, including ARIMA (Makridakis & Hibon, 1997) and ETS (Hyndman & Athanasopoulos, 2018), along with the Gradient Boosting Decision Tree (GBDT) model, XGBoost, into the ProbTSframework. The results in Table 9 clearly demonstrate the superior performance of deep learning methods over simple statistical baselines, emphasizing the importance of capturing non-linear dependencies for accurate forecasts. Notably, ARIMA and ETS exhibit varied performance across different data characteristics. ARIMA struggles with datasets like Solar, characterized by weak trending and strong seasonality, while ETS shows better adaptability. Conversely, in cases of strong trending and weak seasonality, as observed in the ’Wikipedia’ dataset, ARIMA significantly outperforms ETS.
Utilizing the implementation from (Elsayed et al., 2021), we find that XGBoost competes well, even surpassing neural network models in certain scenarios. However, for datasets with more complex distributions like ’Solar’ and ’Electricity,’ advanced probabilistic estimation methods demonstrate a substantial advantage over traditional learning methods and point estimation techniques. This highlights the adaptability and strength of advanced probabilistic methods in handling intricate forecasting scenarios.
Table 9: Results of statistical models and GBDT baseline on short-term forecasting datasets.
Model Exchange Rate Solar Electricity Traffic Wikipedia
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
ARIMA $0.009$ $0.009$ $1.000$ $1.000$ $0.164$ $0.164$ $0.461$ $0.461$ $0.348$ $0.348$
ETS $0.011$ $0.011$ $0.580$ $0.580$ $0.121$ $0.121$ $0.413$ $0.413$ $0.685$ $0.685$
ETS-prob $0.008$ $0.011$ $0.795$ $0.695$ $0.123$ $0.129$ $0.380$ $0.433$ $0.625$ $0.697$
XGBoost $0.011$ $0.011$ $0.599$ $0.599$ $0.074$ $0.074$ $0.196$ $0.196$
DLinear $0.012_{.001}$ $0.012_{.001}$ $0.547_{.009}$ $0.547_{.009}$ $0.095_{.006}$ $0.095_{.006}$ $0.273_{.012}$ $0.273_{.012}$ $1.046_{.037}$ $1.046_{.037}$
PatchTST $0.010_{.000}$ $0.010_{.000}$ $0.496_{.002}$ $0.496_{.002}$ $0.076_{.001}$ $0.076_{.001}$ $0.202_{.001}$ $0.202_{.001}$ $0.257_{.001}$ $0.257_{.001}$
TimesNet $0.011_{.001}$ $0.011_{.001}$ $0.507_{.019}$ $0.507_{.019}$ $0.071_{.002}$ $0.071_{.002}$ $0.205_{.002}$ $0.205_{.002}$ $0.304_{.002}$ $0.304_{.002}$
GRU NVP $0.016_{.003}$ $0.020_{.003}$ $0.396_{.021}$ $0.507_{.022}$ $0.055_{.002}$ $0.073_{.003}$ $0.161_{.006}$ $0.203_{.009}$ $0.282_{.003}$ $0.330_{.003}$
GRU MAF $0.015_{.001}$ $0.020_{.001}$ $0.386_{.026}$ $0.492_{.027}$ $0.051_{.001}$ $0.067_{.001}$ $0.131_{.006}$ $0.165_{.009}$ $0.281_{.004}$ $0.337_{.005}$
Trans MAF $0.011_{.001}$ $0.014_{.001}$ $0.400_{.022}$ $0.503_{.022}$ $0.054_{.004}$ $0.071_{.005}$ $0.129_{.004}$ $0.165_{.006}$ $0.289_{.008}$ $0.344_{.008}$
TimeGrad $0.011_{.001}$ $0.014_{.002}$ $0.359_{.011}$ $0.445_{.023}$ $0.052_{.001}$ $0.067_{.001}$ $0.164_{.091}$ $0.201_{.115}$ $0.272_{.008}$ $0.327_{.011}$
CSDI $0.008_{.000}$ $0.011_{.000}$ $0.366_{.005}$ $0.484_{.008}$ $0.050_{.001}$ $0.065_{.001}$ $0.146_{.012}$ $0.176_{.013}$ $0.219_{.006}$ $0.259_{.009}$
B.3 Experiments on Univariate Datasets
In pursuit of a comprehensive analysis spanning univariate and multivariate scenarios, we examined a subset of M4 (Makridakis et al., 2020), M5 (Makridakis et al., 2022), and TOURISM datasets (Athanasopoulos et al., 2011)—crucial datasets for univariate time-series forecasting. Table 10 provides a quantitative assessment of the intrinsic characteristics of these new datasets, focusing on trending strength, seasonality, and data distribution complexity, as detailed in our paper. Notably, these datasets, except for M4-Daily may exhibit fewer seasonal patterns, do not introduce particularly unique characteristics.
Table 10: Quantitative assessment of the intrinsic characteristics of the univariate datasets. The JS Div. denotes Jensen–Shannon divergence, where a lower score indicates closer approximations to a Gaussian distribution.
Dataset M4-Weekly M4-Daily M5 TOURISM-Monthly
Trend $F_{T}$ 0.7677 0.9808 0.3443 0.7979
Seasonality $F_{S}$ 0.3401 0.0467 0.2480 0.6826
JS Div. 0.5106 0.4916 0.6011 0.3291
Table 11 presents experimental results for representative methods, consistent with our initial observations. Probabilistic estimation methods like GRU NVP and TimeGrad excel on datasets with complex distributions (e.g., M4-Weekly and M5), while simpler point forecasting methods such as DLinear and PatchTST perform well on datasets with relatively simple data distribution, like TOURISM-Monthly. Both autoregressive and non-autoregressive decoding schemes show comparable performance in short-term forecasting, as discussed in the main paper.”
Table 11: Results on M4, M5, and TOURISM datasets. We utilize a lookback window of 3H, with ’H’ denoting the forecasting horizon.
Model DLinear PatchTST GRU NVP TimeGrad
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
M4-Weekly 0.081 0.081 0.089 0.089 0.066 0.077 0.055 0.065
M4-Daily 0.034 0.034 0.035 0.035 0.030 0.038 0.026 0.032
M5 0.891 0.891 0.898 0.898 0.679 0.864 - -
TOURISM-Monthly 0.168 0.168 0.136 0.136 0.171 0.223 0.152 0.191
B.4 Experiments on Synthetic Datasets
To enhance the rigor of the insights presented, we employ synthetic datasets created with the GluonTS library [^5] https://ts.gluon.ai/stable/tutorials/data_manipulation/index.html, encompassing a baseline dataset and variants with pronounced trends, strong seasonality, and complex data distribution (see Table 12). Specifically, we generate these datasets by superimposing four components - trend, seasonality, noise, and anomaly - each with adjustable intensity parameters. The seasonality component is defined by period hyper-parameters and intensity coefficients; the trend by slope intensity; the noise by Gaussian distribution sampling with adjustable intensity; and the anomaly by occurrence probability and maximum intensity.
Subsequent experiments on these synthetic datasets (refer to Table 13), using representative models, validate the empirical findings established on other datasets with ProbTS. Key observations include the declining performance of autoregressive decoding models, such as TimeGrad, in the presence of increasing trends, improved performance for models using autoregressive decoding with intensifying seasonality, and the competitive performance of probabilistic methods like CSDI in handling more complex data distributions.
Table 12: Quantitative assessment of intrinsic characteristics for synthetic datasets. The JS Div denotes Jensen–Shannon divergence, where a lower score indicates closer approximations to a Gaussian distribution.
Dataset Normal Strong Trend Strong Seasonality Complex Distribution
Trend $F_{T}$ 0.105 0.554 0.105 0.064
Seasonality $F_{S}$ 0.302 0.302 0.791 0.190
JS Div. 0.261 0.248 0.272 0.469
Table 13: Results on synthetic datasets. The look-back window and forecasting horizon are 30.
Model Normal Strong Trend Strong Seasonality Complex Distribution
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
DLinear $0.013$ $0.013$ $0.001$ $0.001$ $0.014$ $0.014$ $0.301$ $0.301$
PatchTST $0.012$ $0.012$ $0.001$ $0.001$ $0.012$ $0.012$ $0.275$ $0.275$
TimeGrad $0.024$ $0.032$ $0.042$ $0.048$ $0.022$ $0.028$ $0.283$ $0.338$
CSDI $0.013$ $0.014$ $0.010$ $0.007$ $0.020$ $0.027$ $0.269$ $0.301$
B.5 Case Study
To intuitively demonstrate the distinct characteristics of point and probabilistic estimations, a case study was conducted on short-term datasets. Figure 6 illustrates that point estimation yields single-valued, deterministic estimates, in contrast to probabilistic methods, which model continuous data distributions as depicted in Figure 7. This modeling of data distributions captures the uncertainty in forecasts, aiding decision-makers in fields such as weather and finance to make more informed choices. It is also observed that while both methods align well with ground truth values in short-term forecasting datasets, they struggle to accurately capture outliers, particularly noted in the Wikipedia dataset.
Figure 6: Point forecasts from the PatchTST model and the ground-truth value on short-term forecasting datasets.
Figure 7: Forecasting intervals from the TimeGrad model and the ground-truth value on short-term forecasting datasets.
B.6 Model Efficiency
For reference, detailed results regarding memory usage and time efficiency for five representative models on long-term forecasting datasets are provided here. Table 14 displays the computation memory of various models with a forecasting horizon set to 96. Additionally, Table 15 compares the inference time of these models on long-term forecasting datasets, illustrating the impact of changes in the forecasting horizon.
[Computational memory. ] Refer to caption         [Inference time.] Refer to caption
Figure 8: Comparison of computational efficiency. The forecasting horizon is set to 96 for calculating memory usage.
Table 14: Computation memory. The batch size is 1 and the prediction horizon is set to 96.
Metric Dataset DLinear PatchTST LSTM NVP TimeGrad CSDI
NPARAMS (MB) ETTm1 0.075 2.145 1.079 1.233 1.720
Electricity 0.076 2.146 3.680 3.472 1.370
Traffic 0.078 2.149 15.926 8.298 1.390
Weather 0.075 2.145 3.085 0.574 1.721
Exchange 0.075 0.135 1.979 0.488 1.720
Max GPU Mem. (GB) ETTm1 0.002 0.009 0.010 0.012 0.027
Electricity 0.060 0.068 0.129 0.128 1.411
Traffic 0.161 0.168 0.361 0.333 9.102
Weather 0.004 0.012 0.021 0.012 0.070
Exchange 0.002 0.002 0.013 0.008 0.030
Table 15: Comparison of inference time (sec./sample).
Model pred len DLinear PatchTST LSTM NVP TimeGrad CSDI
ETTm1 96 0.0003 ± 0.0000 0.0003 ± 0.0000 0.0352 ± 0.0007 4.1067 ± 0.0504 16.3280 ± 0.0747
192 0.0003 ± 0.0000 0.0003 ± 0.0000 0.0697 ± 0.0020 7.8979 ± 0.0403 25.8378 ± 0.3124
336 0.0003 ± 0.0000 0.0003 ± 0.0000 0.1221 ± 0.0044 13.6197 ± 0.1023 39.8832 ± 0.2157
720 0.0004 ± 0.0000 0.0003 ± 0.0000 0.2603 ± 0.0020 28.6074 ± 1.1346 86.1862 ± 0.1863
Electricity 96 0.0004 ± 0.0000 0.0045 ± 0.0001 0.1783 ± 0.0006 13.8439 ± 0.0054 388.3150 ± 0.2155
192 0.0006 ± 0.0000 0.0046 ± 0.0000 0.3700 ± 0.0010 27.6683 ± 0.0368 659.4284 ± 0.2003
336 0.0008 ± 0.0000 0.0049 ± 0.0000 0.7157 ± 0.0028 48.4456 ± 0.0279 -
720 0.0015 ± 0.0000 0.0057 ± 0.0000 2.0785 ± 0.0186 104.1473 ± 0.1465 -
Traffic 96 0.0010 ± 0.0001 0.0102 ± 0.0000 0.3695 ± 0.0022 31.7644 ± 0.0101 -
192 0.0013 ± 0.0000 0.0106 ± 0.0000 0.8287 ± 0.0094 63.5832 ± 0.0060 -
336 0.0020 ± 0.0000 0.0114 ± 0.0001 1.6945 ± 0.0026 111.4147 ± 0.0169 -
720 0.0039 ± 0.0000 0.0137 ± 0.0000 5.0963 ± 0.0018 258.1274 ± 0.6088 -
Weather 96 0.0002 ± 0.0000 0.0004 ± 0.0000 $0$ 0.0016 4.1261 ± 0.0812 37.8984 ± 0.0782
192 0.0003 ± 0.0000 0.0004 ± 0.0000 0.1568 ± 0.0008 8.2913 ± 0.5544 62.0223 ± 0.2329
336 0.0003 ± 0.0000 0.0004 ± 0.0000 0.2482 ± 0.0297 14.2391 ± 0.4891 96.8704 ± 0.2258
720 0.0003 ± 0.0000 0.0005 ± 0.0000 0.5447 ± 0.0249 29.4407 ± 0.3519 216.6044 ± 0.4253
Exchange 96 0.0006 ± 0.0000 0.0004 ± 0.0000 0.0284 ± 0.0001 4.1069 ± 0.0981 17.8655 ± 0.1282
192 0.0007 ± 0.0000 0.0004 ± 0.0000 0.0563 ± 0.0008 8.1576 ± 0.0911 28.5456 ± 0.0873
336 0.0007 ± 0.0000 0.0004 ± 0.0000 0.0966 ± 0.0007 14.4593 ± 0.4466 44.9733 ± 0.3820
720 0.0007 ± 0.0000 0.0004 ± 0.0000 0.2085 ± 0.0046 30.1443 ± 0.5378 97.7417 ± 0.2606
ILI 24 0.0002 ± 0.0000 0.0008 ± 0.0001 0.0080 ± 0.0001 1.0427 ± 0.0190 12.4038 ± 0.1681
192 0.0002 ± 0.0000 0.0008 ± 0.0000 0.0121 ± 0.0003 1.5762 ± 0.0282 12.7187 ± 0.1344
336 0.0002 ± 0.0000 0.0008 ± 0.0000 0.0155 ± 0.0002 2.1344 ± 0.0660 12.7386 ± 0.1868
720 0.0002 ± 0.0000 0.0008 ± 0.0000 0.0196 ± 0.0004 2.5787 ± 0.0594 12.5407 ± 0.0481
B.7 Further Discussion on Cross-channel Interactions
We compile a summary table (Table 16) delineating how models from each branch address the multivariate aspect. Despite a thorough investigation, we have not identified a clear pattern linking the modeling of cross-channel interactions to overall model performance. A notable trend is the prevalent use of a channel-mixing approach in most studies. However, findings are diverse; models like DLinear and PatchTST suggest that processing channels independently can yield superior results, while others like CSDI indicate that explicit modeling of cross-channel interactions offers significant advantages. This diversity underscores the ongoing exploration of the impact of cross-channel interactions on forecasting performance.
Table 16: Summary of how existing models handle multivariate time series.
Model Research branch Process channels independently
Customized neural architectures N-BEATS (Oreshkin et al., 2020)
N-HiTS (Challu et al., 2023)
Autoformer (Wu et al., 2021)
Informer (Zhou et al., 2021)
LTSF-Linear (Zeng et al., 2023) ✗/ ✓
PatchTST (Nie et al., 2023) ✗/ ✓
TimesNet (Wu et al., 2023)
Probabilistic estimation DeepAR (Salinas et al., 2020)
GP-copula (Salinas et al., 2019)
LSTM NVP (Rasul et al., 2021b)
LSTM MAF (Rasul et al., 2021b)
Trans MAF (Rasul et al., 2021b)
TimeGrad (Rasul et al., 2021a)
CSDI (Tashiro et al., 2021)
SPD (Bilos et al., 2023)

### v4

@@
-Position Paper: An Integrated Perspective on Data, Metrics, and Methodology for Deep Time-Series Forecasting
+ProbTS: Benchmarking Point and Distributional Forecasting across Diverse Prediction Horizons
 Abstract
-Deep time-series forecasting plays an integral role in numerous practical applications.
-However, existing research fall short by focusing narrowly on either neural architecture designs for long-term point forecasts or probabilistic models for short-term scenarios.
-By proposing a comprehensive framework, facilitated by a novel tool, ProbTS, that integrates diverse data scenarios, evaluation metrics, and methodological focuses, we aim to transcend the limitations of current forecasting practices.
-Rigorous experimentation uncovers pivotal insights, including the supreme importance of aligning forecasting methodologies with the unique characteristics of the data; the necessity of a broad spectrum of metrics for accurately assessing both point and distributional forecasts; and the challenges inherent in adapting existing forecasting methods to a wider range of scenarios.
-These findings not only challenge conventional approaches but also illuminate promising avenues for future research, suggesting a more nuanced and effective strategy for advancing the field of deep time-series forecasting.
-[^1] The ProbTS tool will be open-sourced.
-time-series forecasting, point forecasting, probabilistic forecasting, neural networks
+Delivering precise point and distributional forecasts across a spectrum of prediction horizons represents a significant and enduring challenge in the application of time-series forecasting within various industries.
+Prior research on developing deep learning models for time-series forecasting has often concentrated on isolated aspects, such as long-term point forecasting or short-term probabilistic estimations.
+This narrow focus may result in skewed methodological choices and hinder the adaptability of these models to uncharted scenarios.
+While there is a rising trend in developing universal forecasting models, a thorough understanding of their advantages and drawbacks, especially regarding essential forecasting needs like point and distributional forecasts across short and long horizons, is still lacking.
+In this paper, we present ProbTS, a benchmark tool designed as a unified platform to evaluate these fundamental forecasting needs and to conduct a rigorous comparative analysis of numerous cutting-edge studies from recent years.
+We dissect the distinctive data characteristics arising from disparate forecasting requirements and elucidate how these characteristics can skew methodological preferences in typical research trajectories, which often fail to fully accommodate essential forecasting needs.
+Building on this, we examine the latest models for universal time-series forecasting and discover that our analyses of methodological strengths and weaknesses are also applicable to these universal models.
+Finally, we outline the limitations inherent in current research and underscore several avenues for future exploration.
+[^1] Project repository: https://github.com/microsoft/ProbTS
 1 Introduction
-Deep time-series forecasting serves a crucial role in a myriad of practical applications, from traffic flow forecasting (Lv et al., 2014) and renewable energy forecasting (Wang et al., 2019), to diverse forecasting demands in retail (Böse et al., 2017), finance (Hou et al., 2021), and climate (Mudelsee, 2019).
-Each of these domains presents unique data characteristics, such as trend movements (Taylor & Letham, 2018), periodic patterns (Smyl, 2020), complicated mixtures of global seasonality and local variations (Fan et al., 2022), as well as intricate distributional dependencies (Rasul et al., 2021b).
-These data characteristics collectively pose substantial challenges for accurate forecasting, and thereby necessitate the development of sophisticated modeling designs.
-In addition to accurate point forecasts, modern decision-making processes often require robust distributional forecasts to effectively manage uncertainty (Gneiting & Katzfuss, 2014; Hyndman & Athanasopoulos, 2018).
-These collective requirements emphasize the significance for methodologies that are not only versatile across various forecasting scenarios but also capable of delivering both point and distributional forecasts with high fidelity.
-However, these practical needs also highlight a distinct gap in existing research for deep time-series forecasting, which tends to bifurcate into two streams: one focusing on neural architecture designs optimized for point forecasting in long-term scenarios, and the other on advancing probabilistic forecasting, primarily in short-term situations.
-Particularly, one stream of research, exemplified by studies such as (Zhou et al., 2021; Wu et al., 2021; Liu et al., 2022; Zhang et al., 2022; Challu et al., 2023; Wu et al., 2023; Nie et al., 2023), primarily explores neural architecture designs tailored for time-series forecasting.
-These studies, however, mostly focused on point forecasts, and their architectural designs are largely driven by long-term forecasting scenarios that exhibit strong trending and seasonal patterns.
-Therefore, it remains unclear how these advancements in neural architecture designs can be effectively extended to distributional forecasts, and whether their specialized designs, which are closely tied to certain data patterns, maintain their effectiveness across varied data scenarios.
-On the other hand, another line of research, including works like (Rasul et al., 2021b, a; Tashiro et al., 2021; Bilos et al., 2023; Kollovieh et al., 2023), emphasizes the adaption of deep generative modeling (Dinh et al., 2017; Papamakarios et al., 2017; Ho et al., 2020) to probabilistic time-series forecasting.
-Despite significant progress in characterizing complex data distributions without the need for a predefined closed-form prior distribution such as Gaussian (Salinas et al., 2019), these probabilistic models have primarily been developed and evaluated in short-term forecasting scenarios.
-Hence, it remains to be seen whether these designs can be effectively transitioned to long-term forecasting, and whether improvements in distributional forecasts also yield benefits in point forecasts, especially when compared to the tailored architectural designs mentioned earlier.
-In a bid to systematically bridge this research gap, in this paper, we propose a holistic framework that integrates diverse data scenarios, comprehensive evaluation metrics, and methodological innovations.
-This approach enables a nuanced analysis of why certain forecasting methodologies outperform others in specific contexts and how they could be optimized to improve both point and distributional forecasting.
-To be more concrete, firstly, in terms of data, we encompass various data scenarios, spanning different domains and forecasting horizons.
-More significantly, we aim to quantify key data characteristics, elucidating why certain methodological choices converge or diverge.
-Secondly, concerning metrics, we adopt a comprehensive set of evaluation strategies, assessing the performance of both point and distributional forecasts.
-Lastly, we revisit a selection of representative studies from different research branches, affirming their advancements in specific methodological focuses and examining their performance in seldom evaluated scenarios.
-Through this integrated lens, we aspire to illuminate previously unaddressed questions, expose overlooked limitations of existing methods in certain situations, and reveal our novel findings that could potentially open up new avenues for future research.
-To achieve these ends, we develop a tool, ProbTS, which incorporates customized data, evaluation, and model modules in alignment with the integrated perspective.
-Rigorous experimentation with this tool has led us to a number of intriguing insights, which are concisely presented below and explored in greater detail in Section 4.
-Firstly, regarding the data, we have discovered that the distinctive characteristics of data in different forecasting scenarios primarily dictate the selection of methodology.
-We have identified several fundamental indicators that assist in distinguishing the distinct needs of diverse forecasting scenarios, such as data from various domains, short-term versus long-term forecasting, and point versus distributional forecasting.
-Notably, these indicators also facilitate our understanding of why certain methodological choices align or diverge in particular cases.
-Secondly, concerning metrics, we have discerned that the use of a comprehensive suite of metrics is critical for both point and distributional forecasts.
-This conclusion stems from our observation that optimal performance in distributional metrics does not necessarily translate to optimal performance in point forecasts, and vice versa.
-Importantly, this suggests that focusing solely on optimizing one type of evaluation objective could result in sub-optimal performance according to other evaluation criteria.
-Thirdly, in relation to the first research branch primarily dedicated to developing new architectures tailored for long-term point forecasting scenarios, we have recognized significant challenges and unresolved questions when adapting them to new scenarios, particularly those with distributional or short-term requirements.
-The integration of these architectures with advanced distribution estimation capabilities remains an open issue.
-Furthermore, the advantages of specific designs created for long-term scenarios diminish significantly in short-term cases.
-Another interesting observation is that all these methods consistently favor non-autoregressive decoding schemes over their autoregressive counterparts for multi-step forecasting, likely due to the pronounced trends in many long-term forecasting scenarios.
-In terms of the second research branch, which focuses on developing advanced probabilistic forecasting capabilities, primarily in short-term scenarios, we have observed that extending existing probabilistic models to long-term scenarios presents significant challenges.
-Specifically, they often underperform those point forecasting methods even in producing distributional forecasts.
-Different from the first branch, these probabilistic methods exhibit a balanced preference for autoregressive and non-autoregressive decoding schemes.
-Our analyses reveal that autoregressive probabilistic methods tend to propagate more errors as the forecasting horizon increases, and non-autoregressive probabilistic methods demand substantial computational resources, both in terms of memory and time, as the forecasting horizon expands significantly.
-Interestingly, we also discovered that autoregressive probabilistic methods exhibit significant advantages in dealing with strong seasonality, even in long-term forecasting scenarios.
-This suggests that autoregressive decoding holds substantial potential for long-term forecasting, provided we can effectively mitigate the error-propagation effect amplified by pronounced trending.
-Lastly, by leveraging our versatile tool that facilitates a deeper understanding of the interplay between data characteristics, forecasting performance, and methodological choices, we underscore several under-explored yet valuable directions that not only illuminate the path forward for academic research but also hold profound implications for practical applications in a multitude of sectors.
+Time-series forecasting has extensive applications in various industries, including traffic flow forecasting [42], renewable energy forecasting [65], and diverse forecasting demands in retail [9], finance [30], and climate [47].
+It is crucial to provide forecasts across different prediction horizons, addressing both short- and long-term planning needs [14, 27, 5, 59].
+Moreover, modern decision-making processes typically require not only point forecasts to quantify planning efficiency but also robust distributional estimations to manage uncertainty effectively [25, 31].
+The fundamental need to produce accurate point and distributional forecasts across various horizons presents significant challenges to existing forecasting approaches.
+Nevertheless, much of the previous research on developing deep learning models for time-series forecasting has often focused on isolated aspects, such as long-term point forecasting or short-term distribution estimations.
+This narrow focus may result in skewed methodological choices and hinder the adaptability of these models to rarely evaluated scenarios.
+For example, studies such as [74, 70, 38, 73, 12, 69, 48, 71, 39] have primarily explored neural architecture designs tailored for long-term point forecasting with strong trending and seasonal patterns.
+However, it remains unclear how these advancements can be effectively extended to capture complicated distributions and whether these designs maintain their effectiveness in short-term scenarios.
+Conversely, research such as [56, 55, 60, 8, 33] adapts deep generative models [18, 29] for probabilistic forecasting, specializing in characterizing complex data distributions.
+Yet, these models have mainly been developed and evaluated in short-term scenarios, leaving questions about their effectiveness in long-term forecasting and their ability to preserve point forecasting performance.
+Despite the recent surge in building time-series foundation models over the past year [19, 54, 15, 11, 16, 26, 40, 21, 68, 3, 24], our understanding of their advantages and limitations, especially regarding essential forecasting needs like point and distributional forecasts across various horizons, is still limited.
+Many of these models claim to support arbitrary prediction horizons, employing different mechanisms that come with their own set of advantages and drawbacks.
+Among them, a select few offer capabilities for distributional forecasting, which, however, are typically confined to predefined closed-form distributions [54, 68] or discrete distributions with value quantization [3].
+The emergence of these foundation models has brought about unprecedented zero-shot forecasting capabilities.
+Consequently, it is both timely and crucial to delve into an evaluation of their strengths and weaknesses, especially in relation to the fundamental forecasting needs mentioned earlier.
+In this study, we present ProbTS, a benchmark tool crafted to serve as a comprehensive platform for assessing those key forecasting needs and for performing a detailed comparison of several state-of-the-art models developed in recent years.
+To address the core forecasting requirements, ProbTS includes a broad array of datasets and spans various forecasting horizons.
+It also utilizes both point and distributional metrics to facilitate a thorough performance evaluation.
+Our research reveals that the specific data characteristics inherent to different forecasting requirements often play a crucial role in guiding the selection of model designs.
+As a result, it is crucial to have a comprehensive view of the essential forecasting needs.
+To aid in the analysis and interpretation of performance, we measure three essential data characteristics in ProbTS: the strength of trends and seasonality, and the complexity of the data distribution.
+Moreover, we have explicitly distinguished two fundamental methodological aspects within ProbTS that differentiate the existing forecasting models, largely influencing their pros and cons.
+The first aspect pertains to the approach for enabling distributional forecasts.
+This ranges from models focused solely on point forecasting [48, 39], to those employing pre-defined distribution heads based on certain assumptions about the data distribution [54, 68], and to sophisticated neural probabilistic modules designed to flexibly adapt to various data distributions [55, 60].
+The second aspect is the decoding scheme used to generate multi-step forecasts, which can be either autoregressive (AR) or non-autoregressive (NAR).
+By utilizing ProbTS, we conduct a systematic comparison between studies that focus on long-term point forecasting and those aimed at short-term distributional estimation, employing various forecasting horizons and evaluation metrics.
+Our overarching finding is that the strengths of these methods tend to diminish in scenarios they are rarely evaluated in, highlighting several important but unresolved research questions.
+Notably, while recent probabilistic forecasting approaches have shown proficiency in short-term distribution estimation, we discover that long-term distributional forecasting remains a significant challenge.
+This challenge stems from the difficulty in achieving adaptive distribution estimation that remains both efficient and effective as the prediction horizon extends—a topic that has not been thoroughly investigated in existing literature.
+Additionally, our analysis uncovers a clear divide in the choice of decoding schemes: most long-term point forecasting methods opt for NAR, whereas choices in short-term forecasting studies are more evenly split.
+Further investigation suggests that this preference may be linked to the challenge of error accumulation in AR decoding, particularly as the forecasting horizon lengthens or when dealing with strong trends.
+However, we also discover that AR decoding tends to perform better in scenarios with pronounced seasonality, indicating a potential area for further exploration in decoding strategies, especially for long-term forecasting.
+Together with the notable inefficiency of existing NAR-based probabilistic methods, such as CSDI [60], for long-term forecasts, our comparison between NAR and AR decoding encourages a wider exploration of decoding strategies in future research.
+Furthermore, we have expanded the analytical framework of ProbTS to include an examination of several very recently developed time-series foundation models, which has allowed us to re-validate some of our earlier findings.
+Interestingly, there appears to be a relatively even split in their preference for AR and NAR decoding schemes.
+Our analysis reaffirms the limitation of AR in handling time-series data, as we observe that AR-based foundation models tend to excel at shorter horizons.
+However, their performance advantages often significantly diminish over longer forecasting periods.
+This underscores the critical need for future research to focus on addressing the issue of error accumulation in AR-based foundation models.
+Besides, our exploration reveals that current probabilistic foundation models may face challenges when dealing with complex data distributions.
+This observation suggests that the integration of more sophisticated distribution estimation techniques could enhance the development of time-series foundation models.
+In summary, we have made the following contributions.
+  • Introduction of ProbTS, a benchmark tool designed for a thorough evaluation of essential forecasting needs, towards precise point and probabilistic forecasting across varied horizons.
+  • Comprehensive analysis of methodological variations within forecasting models, especially regarding distributional estimation methods and decoding schemes (AR vs.
+NAR), which illuminates significant yet previously underexplored research challenges.
+  • Extension of our analytical framework to include the latest time-series foundation models, providing insights into the implications of their methodological choices and underscoring important directions for future research endeavors.
 2 Related Work
-As outlined in the introduction, deep time-series forecasting can be divided into two primary branches.
-This section provides a more thorough review of these branches, as well as a brief overview of existing forecasting toolkits.
-Besides, we include a refined comparison in Appendix A.1.
-Neural Architecture Designs in Time-series Forecasting.
-Significant research efforts have been devoted to improving neural architecture designs for time-series forecasting, with contributions ranging from extensions of multi-layer perceptrons (Oreshkin et al., 2020; Zhang et al., 2022) to customized recurrent or convolutional neural networks (Lai et al., 2018; LIU et al., 2022), and Transformer-based variants (Vaswani et al., 2017; Zhou et al., 2022).
-These studies typically employ non-autoregressive decoding schemes and focus on long-term forecasting scenarios, often characterized by strong trend and seasonality patterns.
-However, they predominantly limit themselves to point forecasts, capturing only the average future variations.
-While approaches like quantile regression (Wen et al., 2017; Lim et al., 2021) can mitigate this limitation, they still cannot replace the inherent capture of the data distribution.
-Probabilistic Estimation in Time-series Forecasting.
-In contrast, the second branch of research, known as deep probabilistic time-series forecasting, focuses on leveraging deep neural networks to capture the complex data distribution of future time series.
-These studies span from early approaches using pre-defined likelihood functions (Rangapuram et al., 2018; Salinas et al., 2020), to Gaussian copulas (Salinas et al., 2019; Drouin et al., 2022), and to more recent methods exploiting advanced deep generative models (Rasul et al., 2021b; Bilos et al., 2023).
-While providing probabilistic forecasts, they mainly evaluate short-term forecasting scenarios.
-Interestingly, unlike the first branch, this branch utilizes both autoregressive (Rasul et al., 2021b, a) and non-autoregressive (Tashiro et al., 2021; Bilos et al., 2023; Kollovieh et al., 2023) decoding schemes.
-Furthermore, this branch tends to use general neural architectures instead of customizing them.
-While a few methods (Tashiro et al., 2021; Li et al., 2022; Bergsma et al., 2022) incorporate customized network designs, more studies (de Bézenac et al., 2020; Rasul et al., 2021b, a; Bilos et al., 2023; Drouin et al., 2022) leverage standardized neural networks to encode time-series representations.
+Classical Time-series Forecasting Models
+In recent years, classical research in time-series forecasting has bifurcated into two distinct but complementary streams.
+The first stream has concentrated on refining neural architecture designs for long-term forecasting, primarily employing non-autoregressive decoding schemes to address scenarios with pronounced trend and seasonality.
+This stream has evolved from enhancing multi-layer perceptrons [52, 73] to developing specialized recurrent or convolutional neural networks [34, 37], and introducing Transformer-based models [64, 48, 39].
+Despite achieving advancements in point forecasts, these efforts mainly capture average future changes, with only a few adopting approaches like quantile regression to partially overcome this limitation [67, 36].
+On the other hand, the second stream, probabilistic time-series forecasting specializes in capturing the intricate data distribution of future time series.
+It encompasses a spectrum of techniques, from utilizing predefined likelihood functions [53, 58] and Gaussian copulas [57, 20] to exploring advanced deep generative models [56, 8].
+Unlike the first stream, this branch employs both AR [56, 55] and NAR decoding schemes [60, 8, 33], often utilizing standard neural network architectures to represent time series [17, 56, 55, 8, 20], though some studies propose customized designs [60, 35, 6].
+Together, these streams highlight the diverse approaches to forecasting, ranging from point predictions focusing on the mean future variations to probabilistic forecasts that capture the full distribution of future values.
+In Appendix A.1, we summarize a comparison of these models on the coverage of essential forecasting needs and their methodological preferences.
+Universal Time-series Foundation Models
+Over the past year, the development of time-series foundation models has greatly accelerated, driven by the success of language foundation models [10].
+This wave has seen models such as Lag-Llama [54], TimesFM [16], Timer [40], and Chronos [3] adopting the decoder-only Transformer architecture with an AR decoding scheme.
+Conversely, models like ForecastPFN [19], MOIRAI [68], TTM [21], and UniTS [24] employ the NAR decoding, often using variable-length placeholders to indicate prediction positions for different horizons.
+Probabilistic forecasting is less common, with MOIRAI and Lag-Llama integrating pre-defined distribution heads (Student-t for Lag-Llama and a mixture for MOIRAI) while Chronos uses quantized bins to accommodate time-series values and adopts Softmax outputs for distribution approximation.
+The strategic choice between AR and NAR decoding and the method for distributional estimation highlight distinct trade-offs.
+For an extensive comparison, see Appendix A.2.
 Toolkits for Time-series Forecasting.
 We observe a plethora of toolkits that have been developed for time-series forecasting.
-These range from those primarily designed for point forecasting, such as Prophet (Taylor & Letham, 2018), sktime (Löning et al., 2019), tsai (Oguiza, 2022), and TSlib (Wu et al., 2023), to others that incorporate probabilistic forecasting, including GluonTS (Alexandrov et al., 2020), PyTorchTS (Rasul et al., 2021b), PyTorchForecasting222github.com/jdb78/pytorch-forecasting, and NeuralForecast333github.com/Nixtla/neuralforecast.
-In the development of ProbTS, we drew upon these existing tools, particularly PyTorchTS and GluonTS.
-However, ProbTS distinguishes itself by its specialized design towards the integrated perspective proposed in this paper, encompassing modularized designs across data, methodology, and evaluation components.
-Furthermore, we incorporate more state-of-the-art forecasting models in ProbTS, by migrating their official implementations and conducting extensive hyper-parameter tuning to ensure optimal performance.
+These range from those primarily designed for point forecasting, such as Prophet [61], sktime [41], tsai [51], and TSlib [69], to others that incorporate probabilistic forecasting, including GluonTS [2], PyTorchTS [56], PyTorchForecasting222github.com/jdb78/pytorch-forecasting, and NeuralForecast333github.com/Nixtla/neuralforecast.
+In creating ProbTS, we built upon the foundations laid by tools like PyTorchTS, GluonTS, and TSlib.
+Our unique contribution is a detailed approach that supports both precise point and probabilistic forecasting over various horizons, and examines methodological differences in forecasting models, especially regarding distributional estimation and decoding schemes (AR vs.
+NAR).
+Additionally, ProbTS integrates cutting-edge time-series foundation models, making it a comprehensive benchmark tool for tackling current and future challenges in time-series forecasting.
+A comparison of ProbTS with existing toolkits, focusing on functionalities and features, is provided in Appendix A.3.
 3 The ProbTS Tool
@@
 Figure 1: An overview of ProbTS.
-Data.
+Data
 We aggregate publicly accessible datasets used for both short-term and long-term forecasting.
@@
 Consequently, we have developed quantified indicators, such as trend and seasonality strengths, along with non-Gaussianity to indicate the complexity of data distribution within a forecasting window.
-Detailed information about dataset statistics, visualization analyses, and quantified measures can be found in Appendix A.2A.3A.4, and A.5.
+Detailed information about dataset statistics, visualization analyses, and quantified measures can be found in Appendix B.1.1B.1.2B.1.3, and B.1.4.
 The quantified measurements for all forecasting scenarios are compiled in Table 1.
-Metrics.
+Metrics
 ProbTS incorporates a broad range of evaluation metrics to enable a thorough assessment of both point and distributional forecasts.
-These metrics are elaborated in detail in Appendix A.6.
+These metrics are elaborated in detail in Appendix B.2.
 In this paper, we primarily use the normalized mean absolute error (NMAE) for point forecasts and the continuous ranked probability score (CRPS) for distributional forecasts to succinctly communicate the critical insights discovered.
-It is noteworthy that some methods reproduced in ProbTS, their original papers reported certain point forecast metrics before de-normalizing forecasts to the initial scale (Zeng et al., 2023; Wu et al., 2023; Nie et al., 2023) or primarily reveal aggregated distributional metrics over all time-series variates, namely CRPS-sum (Salinas et al., 2019; Rasul et al., 2021a, b).
-We have diligently ensured our reproduced results align with their reported results and utilized the unified metrics in this study to offer a comprehensive and fair comparison.
-Model.
-The design of our model module is modularized to accommodate various methodological focuses, such as point versus distributional forecasts, customized versus general neural architectures for encoding time-series representations, and autoregressive versus non-autoregressive decoding schemes.
-This design enables us to easily integrate with existing methods by selecting different routes and configuring specific components appropriately.
-Specifically, our design is grounded on the following formulation for time-series forecasting that considers different forecasting paradigms and decoding schemes.
-Formulation.
+It is noteworthy that some methods reproduced in ProbTS, their original papers reported certain point forecast metrics before de-normalizing forecasts to the initial scale [72, 69, 48] or primarily reveal aggregated distributional metrics over all time-series variates, namely CRPS-sum [57, 55, 56].
+We have verified our reproduced results align with their reported results and utilized the unified metrics in this study to offer a comprehensive and fair comparison of these studies from different research threads.
+Model
+The model module in ProbTS explicitly differentiates critical methodological decisions, especially the decoding scheme (AR vs NAR) and the distributional estimation approach.
+Specifically, we employ the following mathematical formulation.
 We denote an element of a multivariate time series as $x^{k}_{t}\in\mathbb{R}$, where $k$ represents the variate index and $t$ denotes the time index.
@@
 Each $x^{k}_{t}$ is associated with covariates $\bm{c}_{t}^{k}\in\mathbb{R}^{N}$, which encapsulates auxiliary information about the observations.
-Given a length-$T$ forecast horizon a length-$L$ observation history $\bm{x}_{t-L:t}$ and corresponding covariates $\bm{c}_{t-L:t}$, the objective in time series forecasting is to generate the vector of future values $\bm{x}_{t+1:t+T}$.
-In ProbTS, we decouple a model into an encoder $f_{\phi}$ and a forecaster $p_{\theta}$.
+Given a length-$T$ forecast horizon, a length-$L$ observation history $\bm{x}_{t-L:t}$ and corresponding covariates $\bm{c}_{t-L:t}$, the objective in time series forecasting is to generate the vector of future values $\bm{x}_{t+1:t+T}$.
+To represent point and distributional forecasting in a unified way, here we divide a model into an encoder $f_{\phi}$ and a forecaster $p_{\theta}$.
 An encoder is tasked with generating expressive hidden states $\bm{h}\in\mathbb{R}^{D}$.
@@
 A forecaster $p_{\theta}$ is employed either to directly estimate point forecasts as $\hat{\bm{x}}_{t}=p_{\theta}(\bm{h}_{t})$, or to perform sampling based on the estimated probabilistic distributions as $\hat{\bm{x}}_{t}\sim p_{\theta}(\bm{x}_{t}|\bm{h}_{t})$.
-Table 1: This table presents a quantitative assessment of the inherent characteristics for all forecasting scenarios, each corresponding to a dataset with a specific forecasting horizon.
-We use the suffixes ”-S” and ”-L” to differentiate between short-term and long-term scenarios.
+Table 1: This table includes a quantitative assessment of the inherent characteristics for all forecasting scenarios, each corresponding to a dataset with a specific forecasting horizon.
+We use the suffixes "-S" and "-L" to differentiate between short-term and long-term scenarios.
 Quantified indicators encompass trend and seasonality strengths, as well as non-Gaussianity, where a higher value signifies a greater deviation from a Gaussian distribution.
 Dataset-Horizon Exchange-S Solar-S Electricity-S Traffic-S Wikipedia-S ETTm1-L ETTm2-L
-Trend 0.9982 0.1688 0.6443 0.2880 0.5253 0.9462 0.9770
-Seasonality 0.1256 0.8592 0.8323 0.6656 0.2234 0.0105 0.0612
+Trend $F_{T}$ 0.9982 0.1688 0.6443 0.2880 0.5253 0.9462 0.9770
+Seasonality $F_{S}$ 0.1256 0.8592 0.8323 0.6656 0.2234 0.0105 0.0612
 Non-Gaussianity 0.2967 0.5004 0.3579 0.2991 0.2751 0.0833 0.1701
 Dataset-Horizon ETTh1-L ETTh2-L Electricity-L Traffic-L Weather-L Exchange-L ILI-L
-Trend 0.7728 0.9412 0.6476 0.1632 0.9612 0.9978 0.5438
+Trend $F_{T}$ 0.7728 0.9412 0.6476 0.1632 0.9612 0.9978 0.5438
 Seasonality $F_{S}$ 0.4772 0.3608 0.8344 0.6798 0.2657 0.1349 0.6075
 Non-Gaussianity 0.0719 0.1422 0.1533 0.1378 0.1727 0.1082 0.1112
-4 Analyses
-Table 2: Results ($\textrm{mean}_{\textrm{std}}$) on short-term forecasting scenarios, each containing five independent runs with different seeds.
-Model Exchange Rate Solar Electricity Traffic Wikipedia
-CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
-Glob. mean $0.188$ $0.188$ $1.403$ $1.403$ $0.412$ $0.412$ $0.540$ $0.540$ $0.577$ $0.577$
-Batch mean $0.012$ $0.012$ $1.244$ $1.244$ $0.365$ $0.365$ $0.503$ $0.503$ $0.336$ $0.336$
-Linear $0.012_{.001}$ $0.012_{.001}$ $0.704_{.036}$ $0.704_{.036}$ $0.138_{.009}$ $0.138_{.009}$ $0.327_{.032}$ $0.327_{.032}$ $0.874_{.151}$ $0.874_{.151}$
-GRU $0.013_{.002}$ $0.013_{.002}$ $0.594_{.144}$ $0.594_{.144}$ $0.134_{.009}$ $0.134_{.009}$ $0.193_{.002}$ $0.193_{.002}$ $0.394_{.013}$ $0.394_{.013}$
-Transformer $0.016_{.001}$ $0.016_{.001}$ $0.538_{.066}$ $0.538_{.066}$ $0.115_{.005}$ $0.115_{.005}$ $0.204_{.006}$ $0.204_{.006}$ $0.408_{.011}$ $0.408_{.011}$
-N-HiTS $0.012_{.000}$ $0.012_{.000}$ $0.572_{.020}$ $0.572_{.020}$ $0.074_{.003}$ $0.074_{.003}$ $0.193_{.002}$ $0.193_{.002}$ $0.332_{.011}$ $0.332_{.011}$
-NLinear $0.010_{.000}$ $0.010_{.000}$ $0.560_{.002}$ $0.560_{.002}$ $0.083_{.002}$ $0.083_{.002}$ $0.233_{.001}$ $0.233_{.001}$ $0.321_{.001}$ $0.321_{.001}$
-DLinear $0.012_{.001}$ $0.012_{.001}$ $0.547_{.009}$ $0.547_{.009}$ $0.076_{.003}$ $0.076_{.003}$ $0.250_{.002}$ $0.250_{.002}$ $0.412_{.001}$ $0.412_{.001}$
-PatchTST $0.010_{.000}$ $0.010_{.000}$ $0.496_{.002}$ $0.496_{.002}$ $0.067_{.001}$ $0.067_{.001}$ $0.202_{.001}$ $0.202_{.001}$ $0.257_{.001}$ $0.257_{.001}$
-TimesNet $0.011_{.001}$ $0.011_{.001}$ $0.507_{.019}$ $0.507_{.019}$ $0.071_{.002}$ $0.071_{.002}$ $0.205_{.002}$ $0.205_{.002}$ $0.304_{.002}$ $0.304_{.002}$
-GRU NVP $0.016_{.003}$ $0.020_{.003}$ $0.396_{.021}$ $0.507_{.022}$ $0.055_{.002}$ $0.073_{.003}$ $0.161_{.006}$ $0.203_{.009}$ $0.282_{.003}$ $0.330_{.003}$
-GRU MAF $0.015_{.001}$ $0.020_{.001}$ $0.386_{.026}$ $0.492_{.027}$ $0.051_{.001}$ $0.067_{.001}$ $0.131_{.006}$ $0.165_{.009}$ $0.281_{.004}$ $0.337_{.005}$
-Trans MAF $0.011_{.001}$ $0.014_{.001}$ $0.400_{.022}$ $0.503_{.022}$ $0.054_{.004}$ $0.071_{.005}$ $0.129_{.004}$ $0.165_{.006}$ $0.289_{.008}$ $0.344_{.008}$
-TimeGrad $0.011_{.001}$ $0.014_{.002}$ $0.359_{.011}$ $0.445_{.023}$ $0.052_{.001}$ $0.067_{.001}$ $0.164_{.091}$ $0.201_{.115}$ $0.272_{.008}$ $0.327_{.011}$
-CSDI $0.008_{.000}$ $0.011_{.000}$ $0.366_{.005}$ $0.484_{.008}$ $0.050_{.001}$ $0.065_{.001}$ $0.146_{.012}$ $0.176_{.013}$ $0.219_{.006}$ $0.259_{.009}$
-Table 3: Results ($\textrm{mean}_{\textrm{std}}$) on long-term forecasting scenarios, each containing five independent runs with different seeds.
-The input sequence length is set to 36 for the ILI dataset and 96 for the others.
-Due to the excessive time and memory consumption of CSDI in producing long-term forecasts, its results are unavailable in some datasets.
-Model pred DLinear PatchTST GRU NVP TimeGrad CSDI
-len CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
-ETTm1 96 $0.282_{.002}$ $0.282_{.002}$ $0.272_{.001}$ $0.272_{.001}$ $0.383_{.053}$ $0.488_{.058}$ $0.522_{.105}$ $0.645_{.129}$ $0.236_{.006}$ $0.308_{.005}$
-192 $0.309_{.004}$ $0.309_{.004}$ $0.295_{.001}$ $0.295_{.001}$ $0.396_{.030}$ $0.514_{.042}$ $0.603_{.092}$ $0.748_{.084}$ $0.291_{.025}$ $0.377_{.026}$
-336 $0.338_{.008}$ $0.338_{.008}$ $0.323_{.001}$ $0.323_{.001}$ $0.486_{.032}$ $0.630_{.029}$ $0.601_{.028}$ $0.759_{.015}$ $0.322_{.033}$ $0.419_{.042}$
-720 $0.387_{.006}$ $0.387_{.006}$ $0.353_{.001}$ $0.353_{.001}$ $0.546_{.036}$ $0.707_{.050}$ $0.621_{.037}$ $0.793_{.034}$ $0.448_{.038}$ $0.578_{.051}$
-ETTm2 96 $0.138_{.000}$ $0.138_{.000}$ $0.132_{.001}$ $0.132_{.001}$ $0.319_{.044}$ $0.413_{.059}$ $0.427_{.042}$ $0.525_{.047}$ $0.115_{.009}$ $0.146_{.012}$
-192 $0.163_{.003}$ $0.163_{.003}$ $0.157_{.001}$ $0.157_{.001}$ $0.326_{.025}$ $0.427_{.033}$ $0.424_{.061}$ $0.530_{.060}$ $0.147_{.008}$ $0.189_{.012}$
-336 $0.188_{.001}$ $0.188_{.001}$ $0.176_{.000}$ $0.176_{.000}$ $0.449_{.145}$ $0.580_{.169}$ $0.469_{.049}$ $0.566_{.047}$ $0.190_{.018}$ $0.248_{.024}$
-720 $0.219_{.003}$ $0.219_{.003}$ $0.205_{.001}$ $0.205_{.001}$ $0.561_{.273}$ $0.749_{.385}$ $0.470_{.054}$ $0.561_{.044}$ $0.239_{.035}$ $0.306_{.040}$
-ETTh1 96 $0.352_{.011}$ $0.352_{.011}$ $0.328_{.003}$ $0.328_{.003}$ $0.379_{.030}$ $0.481_{.037}$ $0.455_{.046}$ $0.585_{.058}$ $0.437_{.018}$ $0.557_{.022}$
-192 $0.393_{.001}$ $0.393_{.001}$ $0.359_{.002}$ $0.359_{.002}$ $0.425_{.019}$ $0.531_{.018}$ $0.516_{.038}$ $0.680_{.058}$ $0.496_{.051}$ $0.625_{.065}$
-336 $0.419_{.007}$ $0.419_{.007}$ $0.384_{.002}$ $0.384_{.002}$ $0.458_{.054}$ $0.580_{.064}$ $0.512_{.026}$ $0.666_{.047}$ $0.454_{.025}$ $0.574_{.026}$
-720 $0.502_{.029}$ $0.502_{.029}$ $0.397_{.002}$ $0.397_{.002}$ $0.502_{.039}$ $0.643_{.046}$ $0.523_{.027}$ $0.672_{.015}$ $0.528_{.012}$ $0.657_{.014}$
-ETTh2 96 $0.211_{.027}$ $0.211_{.027}$ $0.177_{.000}$ $0.177_{.000}$ $0.432_{.141}$ $0.548_{.158}$ $0.358_{.026}$ $0.448_{.031}$ $0.164_{.013}$ $0.214_{.018}$
-192 $0.238_{.028}$ $0.238_{.028}$ $0.201_{.001}$ $0.201_{.001}$ $0.625_{.170}$ $0.766_{.223}$ $0.457_{.081}$ $0.575_{.089}$ $0.226_{.018}$ $0.294_{.027}$
-336 $0.284_{.008}$ $0.284_{.008}$ $0.240_{.001}$ $0.240_{.001}$ $0.793_{.319}$ $0.942_{.408}$ $0.481_{.078}$ $0.606_{.095}$ $0.274_{.022}$ $0.353_{.028}$
-720 $0.307_{.000}$ $0.307_{.000}$ $0.252_{.000}$ $0.252_{.000}$ $0.539_{.090}$ $0.688_{.161}$ $0.445_{.016}$ $0.550_{.018}$ $0.302_{.040}$ $0.382_{.030}$
-Electricity 96 $0.090_{.001}$ $0.090_{.001}$ $0.086_{.001}$ $0.086_{.001}$ $0.094_{.003}$ $0.118_{.003}$ $0.096_{.002}$ $0.119_{.003}$ $0.153_{.137}$ $0.203_{.189}$
-192 $0.095_{.001}$ $0.095_{.001}$ $0.092_{.001}$ $0.092_{.001}$ $0.097_{.002}$ $0.121_{.003}$ $0.100_{.004}$ $0.124_{.005}$ $0.200_{.094}$ $0.264_{.129}$
-336 $0.104_{.000}$ $0.104_{.000}$ $0.100_{.000}$ $0.100_{.000}$ $0.099_{.001}$ $0.123_{.001}$ $0.102_{.007}$ $0.126_{.008}$
-720 $0.122_{.001}$ $0.122_{.001}$ $0.116_{.000}$ $0.116_{.000}$ $0.114_{.013}$ $0.144_{.017}$ $0.108_{.003}$ $0.134_{.004}$
-Traffic 96 $0.356_{.009}$ $0.356_{.009}$ $0.248_{.001}$ $0.248_{.001}$ $0.187_{.002}$ $0.231_{.003}$ $0.202_{.004}$ $0.234_{.006}$
-192 $0.346_{.009}$ $0.346_{.009}$ $0.245_{.001}$ $0.245_{.001}$ $0.192_{.001}$ $0.236_{.002}$ $0.208_{.003}$ $0.239_{.004}$
-336 $0.350_{.008}$ $0.350_{.008}$ $0.257_{.002}$ $0.257_{.002}$ $0.201_{.004}$ $0.248_{.006}$ $0.213_{.003}$ $0.246_{.003}$
-720 $0.365_{.009}$ $0.365_{.009}$ $0.266_{.001}$ $0.266_{.001}$ $0.211_{.004}$ $0.264_{.006}$ $0.220_{.002}$ $0.263_{.001}$
-Weather 96 $0.112_{.001}$ $0.112_{.001}$ $0.087_{.002}$ $0.087_{.002}$ $0.116_{.013}$ $0.145_{.017}$ $0.130_{.017}$ $0.164_{.023}$ $0.068_{.008}$ $0.087_{.012}$
-192 $0.122_{.001}$ $0.122_{.001}$ $0.090_{.001}$ $0.090_{.001}$ $0.122_{.021}$ $0.147_{.025}$ $0.127_{.019}$ $0.158_{.024}$ $0.068_{.006}$ $0.086_{.007}$
-336 $0.130_{.002}$ $0.130_{.002}$ $0.092_{.002}$ $0.092_{.002}$ $0.128_{.011}$ $0.160_{.012}$ $0.130_{.006}$ $0.162_{.006}$ $0.083_{.002}$ $0.098_{.002}$
-720 $0.144_{.001}$ $0.144_{.001}$ $0.094_{.001}$ $0.094_{.001}$ $0.110_{.004}$ $0.135_{.008}$ $0.113_{.011}$ $0.136_{.020}$ $0.087_{.003}$ $0.102_{.005}$
-Exchange 96 $0.024_{.000}$ $0.024_{.000}$ $0.023_{.000}$ $0.023_{.000}$ $0.071_{.006}$ $0.091_{.009}$ $0.068_{.003}$ $0.079_{.002}$ $0.028_{.003}$ $0.036_{.005}$
-192 $0.035_{.000}$ $0.035_{.000}$ $0.034_{.000}$ $0.034_{.000}$ $0.068_{.004}$ $0.087_{.005}$ $0.087_{.013}$ $0.100_{.019}$ $0.045_{.003}$ $0.058_{.005}$
-336 $0.048_{.001}$ $0.048_{.001}$ $0.048_{.000}$ $0.048_{.000}$ $0.072_{.002}$ $0.091_{.002}$ $0.074_{.009}$ $0.086_{.008}$ $0.060_{.004}$ $0.076_{.006}$
-720 $0.075_{.002}$ $0.075_{.002}$ $0.072_{.000}$ $0.072_{.000}$ $0.079_{.009}$ $0.103_{.009}$ $0.099_{.015}$ $0.113_{.016}$ $0.143_{.020}$ $0.173_{.020}$
-ILI 24 $0.213_{.038}$ $0.213_{.038}$ $0.169_{.005}$ $0.169_{.005}$ $0.257_{.003}$ $0.283_{.001}$ $0.275_{.047}$ $0.296_{.044}$ $0.250_{.013}$ $0.263_{.012}$
-36 $0.230_{.015}$ $0.230_{.015}$ $0.156_{.005}$ $0.156_{.005}$ $0.281_{.004}$ $0.307_{.007}$ $0.272_{.057}$ $0.298_{.048}$ $0.285_{.010}$ $0.298_{.011}$
-48 $0.221_{.009}$ $0.221_{.009}$ $0.156_{.008}$ $0.156_{.008}$ $0.288_{.008}$ $0.314_{.009}$ $0.295_{.033}$ $0.320_{.025}$ $0.285_{.036}$ $0.301_{.034}$
-60 $0.230_{.013}$ $0.230_{.013}$ $0.147_{.003}$ $0.147_{.003}$ $0.307_{.005}$ $0.333_{.005}$ $0.295_{.083}$ $0.325_{.068}$ $0.283_{.012}$ $0.299_{.013}$
-[Long-term forecasting (NMAE).] Refer to caption [Long-term forecasting (CRPS).] Refer to caption     Refer to caption
-[Short-term forecasting (NMAE).] Refer to caption [Short-term forecasting (CRPS).] Refer to caption     Refer to caption
-Figure 2: We present plots of the metrics against non-Gaussianity for various models across all forecasting scenarios.
-In each subplot, a vertical line of points represents one forecasting scenario, denoted by a sequence number at the top.
-We distinguish between point and probabilistic forecasting methods using different colors, and we use unique markers to indicate various models.
+4 Results and Analyses
+[Long-term Forecasting (NMAE).] Refer to caption [Long-term Forecasting (CRPS).] Refer to caption
+[Short-term Forecasting (NMAE).] Refer to caption [Short-term Forecasting (CRPS).] Refer to caption
+Figure 2: We present a comprehensive comparison between classical models designed for long-term point forecasting and short-term distributional forecasting across various prediction horizons.
+It utilizes a non-Gaussianity score to highlight the complexity of the data distribution across different datasets.
+The aggregated performance metrics are derived from Tables 9 and 10.
 [CRPS w.r.t. forecasting horizon.] Refer to caption [CRPS w.r.t. trend strength.] Refer to caption
-[CRPS w.r.t. seasonality strength.] Refer to caption     [The CRPS Gap, computed as TimeGrad.CRPS - CSDI.CRPS, w.r.t. trend and seasonal strengths.] Refer to caption
-Figure 3: We analyze CRPS in relation to various factors to scrutinize the strengths and weaknesses of autoregressive probabilistic methods.
-Subplot (a), based on ETTm and ILI datasets, illustrates the error amplification that occurs with increased prediction length.
-Subplots (b) and (c) encompass all long-term forecasting scenarios, aiming to explore the relationship between CRPS and the strengths of trends and seasonality, respectively.
-Subplot (d) includes all forecasting scenarios, with a lighter and smaller circle denoting a preference for autoregressive methods over non-autoregressive ones.
-Utilizing ProbTS, we benchmarked selected methods in both short-term and long-term forecasting scenarios.
-We evaluated these methods using both point and distributional forecast metrics.
-To ensure the reliability of our results and minimize randomness, we performed each experiment five times with different seeds for parameter initialization.
-The results for short-term and long-term forecasting are presented in Table 2 and Table 3, respectively.
-Selected Methods for Comparison.
-We have chosen representative methods from various research branches for comparison.
-Our selection is based on a comprehensive consideration of factors such as state-of-the-art performance, reproducibility, as well as the simplicity, universality, and novelty of the method designs.
-Adhering to these principles, we have selected PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023), N-HiTS (Challu et al., 2023), LTSF-Linear (NLinear and DLinear) (Zeng et al., 2023) from the research line that concentrates on developing customized neural architectures for long-term point forecasting.
-In addition, we have chosen advanced probabilistic forecasting methods for comparison, including those built upon normalizing flows such as GRU NVP, GRU MAF, and Trans MAF (Rasul et al., 2021b), and diffusion models such as TimeGrad (Rasul et al., 2021a) and CSDI (Tashiro et al., 2021).
-We also include general network architectures, such as Linear, GRU (Chung et al., 2014), and Transformer (Vaswani et al., 2017), to demonstrate the additional benefits brought by advanced architecture designs or probabilistic estimations.
-Simple non-parametric baselines, such as calculating the global mean or batch mean as all forecasts, are also included to assess the predictability of a forecasting scenario.
-4.1 Discrepancies in Point and Distributional Metrics.
-Our analysis revealed that optimal performance in distributional metrics does not guarantee optimal performance in point forecasts.
-This pattern emerged in several instances, notably in Table 3.
-For example, CSDI excels in the CRPS metric but falls short in NMAE across multiple datasets, such as ETTm1, ETTm2, and ETTh2.
-These observations underscore the necessity of a comprehensive evaluation utilizing both NMAE and CRPS metrics to gain a holistic understanding of performance in both point and distributional forecasting.
-It is worth highlighting that these two metrics become equivalent for point forecasting methods when a point forecast is converted into a degenerate distribution, centered on the forecasted point, for CRPS evaluation (Gneiting & Raftery, 2007).
-4.2 Examination of Customized Neural Architectures
-Our findings affirm the progress in this area of research.
-For instance, as observed in Table 3, the state-of-the-art architecture, PatchTST, consistently surpasses other models, especially when assessed using the NMAE metric.
-The only exception is on the Traffic dataset, where PatchTST is outperformed by GRU NVP and TimeGrad.
-This discrepancy is primarily attributable to the different decoding schemes utilized, which will be elaborated in Section 4.3.
-Diminishing Advantages of Customized Architectures in Short-term Forecasting Scenarios.
-Interestingly, the significant superiority of PatchTST diminishes in short-term forecasting scenarios.
-This trend is evident when comparing Table 3 and 2.
-More intuitively, Figures 22 display the relative strengths of point and probabilistic forecasting methods in terms of NMAE, and correlates their NMAE values with our quantified indicator, non-Gaussianity, used for assessing data distribution complexity.
-We further observe that as non-Gaussianity increases, signifying a more complex data distribution, a noticeable performance gap emerges.
-This suggests that the reducing superiority of these customized architectures may be due to their limited capability to handle complex data distributions.
-Challenges in Integrating Advanced Distribution Estimation Capabilities with These Architectures.
-Regrettably, the integration of these architectures with advanced distribution estimation capabilities remains a challenging task.
-As will be discussed subsequently, non-autoregressive probabilistic methods encounter significant efficiency issues in long-term forecasting.
-Intriguingly, all these architectures select a non-autoregressive decoding scheme.
-We will delve into the advantages and disadvantages of autoregressive and non-autoregressive schemes in the next section.
-4.3 Assessment of Probabilistic Forecasting Methods
-At the outset, we validate the efficacy of these probabilistic methods in the short-term forecasting scenarios they were designed for.
-Evidence from Table 2 demonstrates that innovative methods such as CSDI and TimeGrad yield superior results in most cases.
-However, in Long-term Forecasting Scenarios, Probabilistic Methods Fall Short, Even When Evaluated Using Distributional Metrics.
-As depicted in Table 3, in a multitude of long-term forecasting scenarios, customized neural architectures, created specifically with point forecasting objectives in mind, decidedly outperform probabilistic methods, even when evaluated using distributional metrics such as CRPS.
-This suggests that existing probabilistic methods may encounter significant challenges when applied to long-term forecasting, leaving substantial room for improvement in this domain.
-To further illustrate this, we compare the CRPS performance of these models with the non-Gaussianity in both long-term and short-term scenarios, as depicted in Figure 2 and 2, respectively.
-While non-Gaussianity does provide some insights—its values are relatively lower in long-term scenarios compared to short-term contexts—these explanations are not comprehensive.
-This is because it remains unclear to what extent a customized neural model, coupled with effective probabilistic estimations, can perform.
-To delve deeper, we further examine other unique aspects of the existing probabilistic forecasting methods in search of potential clues.
-The Dichotomy of Decoding Scheme Preferences.
-Upon further investigation, we find that these probabilistic forecasting methods exhibit a balanced preference for both autoregressive and non-autoregressive decoding schemes.
-For instance, TimeGrad employs an autoregressive decoding scheme, whereas CSDI utilizes a non-autoregressive decoding approach.
-This contrasts starkly with the aforementioned customized architectures, which solely opt for non-autoregressive decoding.
+[…]
+Figure 3: We explore the challenges faced by current models in conducting long-term distributional forecasting, with insights drawn from Table 10.
+Subplot (a) highlights the substantial error increases in AR-based probabilistic models (e.g., GRU NVP, TimeGrad) when applied to ETTm and ILI datasets.
+Subplots (b) and (c) examine how the strengths of trends and seasonality, respectively, affect the overall performance across all long-term forecasting datasets and horizons.
+Subplot (d) further investigates the combined effects of trend and seasonality, using lighter and smaller circles to indicate situations where AR-based models are favored over NAR-based ones.
+Utilizing ProbTS, we conducted a comprehensive benchmarking and analysis of a diverse range of state-of-the-art models from different strands of research.
+We mainly assessed these models using NAME and CRPS metrics across multiple forecasting horizons, repeating each experiment five times with different seeds to ensure result reliability.
+Selected Models for Comparison.
+Our selection criteria for models focused on a balance of performance, reproducibility, and simplicity.
+For long-term point forecasting, we included models like iTransformer [39], PatchTST [48], TimesNet [69], N-HiTS [12], and LTSF-Linear [72].
+Probabilistic forecasting methods selected include GRU NVP, GRU MAF, Trans MAF [56], TimeGrad [55], and CSDI [60].
+Additionally, general architectures like Linear, GRU [13], and Transformer [64], along with simple non-parametric baselines, were evaluated as a reference.
+For foundation models, reproducible methods such as Lag-Llama [54], TimesFM [16], Timer [40], MOIRAI [68], Chronos [3], and UniTS [24] were included.
+Detailed implementation specifics are in Appendix B.3.
+Due to space constraints, comprehensive comparison results are placed in Appendix C, with detailed results for short-term and long-term forecasting in Tables 9 and 10, respectively.
+Zero-shot evaluations of pre-trained time-series foundation models are detailed in Tables 11 and 12.
+Our evaluation highlights the critical relationship between forecasting requirements, data properties, and modeling strategies.
+It aims to shed light on the strengths and limitations of current approaches, paving the way for uncovering novel research avenues.
+4.1 Analyzing Classical Models for Time-series Forecasting
+We examine traditional non-universal time-series models from distinct research branches: one branch focuses on developing customized neural architectures tailored for long-term point forecasting, while the other branch concentrates on creating advanced probabilistic methods for short-term distributional forecasting.
+Our investigation confirms the effectiveness of these models for their intended purposes.
+However, we observe a notable trend: the strengths of these methods tend to diminish in scenarios where they are seldom tested.
+Diminishing Advantages of Customized Architectures in Short-term Forecasting Scenarios
+The comparative analysis presented in Figures 2 and 2 showcases the performance of point and probabilistic forecasting methods with respect to the NMAE metric.
+These figures also illustrate how NMAE values correlate with non-Gaussianity, a measure we employ to evaluate the complexity of data distributions.
+It becomes evident that customized architectures, originally crafted for long-term forecasting, tend to lose their competitive performance in short-term scenarios.
+This phenomenon could be attributed to the increased importance of accurately characterizing complex data distributions within shorter forecasting windows, where higher non-Gaussianity scores are indicative of this necessity.
+A closer look at Figures 2 and 2 further reveals that the performance disparity measured with CRPS becomes even more pronounced for datasets characterized by significant non-Gaussianity, such as Solar-S.
+This observation underscores the critical need for incorporating short-term patterns and distributional estimation capabilities into the design of new forecasting architectures.
+Significant Performance Degradation for Existing Probabilistic Methods in Long-term Distributional Forecasting
+The performance of current probabilistic forecasting models in long-term scenarios, even when assessed using distributional metrics such as CRPS, reveals notable limitations.
+This is highlighted by the comparison between Figures 2 and 2, which shows a significant drop in performance for models like TimeGrad on ETTm1-L, GRU NVP on ETTh2-L, and Weather-L datasets.
+The decline in performance can be attributed to the fact that these probabilistic models were not specifically designed with the unique challenges of long-term forecasting in mind.
+This oversight has mixed influences.
+On the positive side, the design of these methods has led to a more balanced approach in the choice between AR and NAR decoding schemes, providing a versatile foundation for probabilistic forecasting.
+However, the downside is more significant: no matter using which decoding schemes, existing probabilistic models face considerable challenges when applied to long-term distributional forecasting.
+We will dive deeper into the specific challenges associated with each decoding scheme next.
+Here the performance gap underscores the need for future research to systematically investigate long-term distributional forecasting.
+Different Decoding Schemes & Challenges in Long-term Distributional Forecasting
+Existing probabilistic forecasting methods exhibit a balanced preference for both AR and NAR decoding schemes.
+For instance, TimeGrad employs an AR decoding scheme, whereas CSDI utilizes an NAR approach.
+This contrasts starkly with the aforementioned customized architectures, which solely opt for NAR decoding.
 These two types of decoding schemes, however, confront distinctive challenges when applied to long-term probabilistic forecasting.
-  • Autoregressive probabilistic methods, for example, TimeGrad, are prone to propagate more errors as the forecasting horizon increases or the trend strength enlarges.
-This phenomenon is evident in Figure 3 and 3, where we observe that the performance gap escalates with either extended prediction lengths or intensified trend strengths.
-  • In contrast, non-autoregressive probabilistic methods, like CSDI, struggle with significant efficiency challenges, both in memory usage and time consumption, in long-term forecasting scenarios.
-This issue is further detailed in Appendix Section B.6.
+AR probabilistic models like TimeGrad tend to accumulate errors as they forecast further into the future or as the trend strength increases.
+This issue, which leads to a growing performance gap for longer forecasts or stronger trends, is illustrated in Figures 3 and 3.
+On the other hand, NAR probabilistic models, such as CSDI, face challenges related to memory usage and processing time in long-term forecasts, which are elaborated in Appendix Section D.6.
 Furthermore, as indicated in Table 10, even for relatively small-scale datasets, such as ETTm2 and ETTh1, CSDI’s performance in long-term scenarios is less than optimal.
 This suggests that its learning efficacy is also compromised by the extension of the forecasting horizon.
-The Unexpected Superiority of Autoregressive Decoding Scheme in Addressing Strong Seasonality.
-Interestingly, the autoregressive decoding scheme exhibits a remarkable proficiency in dealing with strong seasonality.
-A prime example is the results for the Traffic dataset shown in Table 3, where both GRU NVP and TimeGrad outperform PatchTST, despite utilizing relatively standard neural architectures without any customization.
-In Figure 3, we further investigate this phenomenon by relating the strength of seasonality to the variance in performance gaps.
-Additionally, in Figure 3, we conduct a collective analysis of the interactions among CRPS Gap, trend, and seasonality strengths, where the CRPS Gap is defined as the difference between the CRPS scores of TimeGrad (autoregressive) and CSDI (non-autoregressive).
-A clear pattern emerges: as the strength of seasonality heightens, autoregressive methods gain more advantages.
-This observation suggests that despite its challenges, the autoregressive decoding scheme should not be dismissed in long-term forecasting.
-Its unique affinity with seasonality, possibly due to superior parameter efficiency, may regain significance if we can develop effective strategies to mitigate the issue of error propagation.
-4.4 Future Directions
-Drawing on our analyses, we propose several promising directions for future research.
-Firstly, the development of hybrid learning objectives that optimize both point and distributional forecasts represents a significant advancement.
-This strategy could foster models capable of delivering more comprehensive predictions, catering to a diverse range of practical applications that necessitate both accurate predictions and well-quantified uncertainties.
-Secondly, there is an urgent need for the effective integration of neural architecture design with advanced probabilistic estimation paradigms.
-Addressing this demand could lay the groundwork for more efficient and precise forecasting models.
-Particularly, overcoming the efficiency challenges associated with non-autoregressive decoding schemes and reducing error propagation in autoregressive probabilistic models, specifically in the presence of marked trends, could substantially enhance the models’ performance and practicality.
-Lastly, the exploration of the necessity and potential designs of customized architectures tailored for short-term forecasting could initiate a novel research field.
-Such exploration necessitates consideration of data characteristics beyond trends and seasonality.
-4.5 Limitations
-Despite the thoroughness of our study, there are several limitations.
-First, due to the vast number of methods available in this field, we could not include all of them in our analysis.
-Our selection is based on a set of criteria, but inevitably, some methods were left out.
-Second, while we have tried to incorporate a diverse range of datasets, there are certainly more datasets beyond those included in our study.
-To mitigate this limitation, we have conducted additional experiments on univariate time-series datasets, which are included in Appendix B.3.
-Finally, our study is primarily empirical in nature.
-While we have made significant efforts to conduct rigorous statistical analyses, we have not provided theoretical proofs for our results.
+The Unexpected Superiority of AR Decoding in Addressing Strong Seasonality
+Despite its drawbacks, AR decoding, as used in TimeGrad, excels in capturing strong seasonality, outperforming models like PatchTST in scenarios such as the Traffic dataset (Table 10).
+This advantage is further analyzed in Figures 3 and 3, showing AR’s increasing benefit with stronger seasonal patterns.
+This suggests AR’s potential in long-term forecasting could be revitalized with solutions to its error accumulation challenge in long horizons.
+[NMAE w.r.t. forecasting horizon.] Refer to caption [$\frac{\text{Model.CRPS}-\text{CSDI.CRPS}}{\text{CSDI.CRPS}}$ w.r.t. non-Gaussianity.] Refer to caption
+Figure 4: We evaluate the efficacy of time-series foundation models for various forecasting horizons and distributional estimation.
+Subplot (a), derived from Table 11 and excluding results from the Electricity dataset, demonstrates the short-term forecasting capabilities and long-term error accumulation of AR-based models.
+Subplot (b), draw from Table 12, investigates short-term distributional estimation, highlighting the performance challenges of foundation models compared to CSDI in handling complex data distributions.
+Note that we include MOIRAI with two different context lengths, 96 and 5000, as context length significantly affects its transfer performance.
+4.2 Analyzing Foundation Models for Universal Time-series Forecasting
+We next explore the capabilities of recent foundation models in universal time-series forecasting, focusing on their performance across different prediction horizons and their ability to estimate distributions, especially regarding their zero-shot transfer capabilities on unseen datasets.
+Table 11 showcases their significant progress, sometimes outperforming traditional models without re-training.
+Using the analytic framework of ProbTS, we delve into their methodological pros and cons.
+Navigating the AR Decoding Challenge over Extended Forecasting Horizons
+Figure 4 illustrates how the performance of various time-series foundation models evolves in relation to expanding forecasting horizons.
+It is evident that for shorter horizons, AR-based foundation models such as TimesFM and Timer exhibit highly competitive performance, on par with NAR-based models like MOIRAI.
+However, the advantage of NAR-based decoding becomes increasingly apparent as the forecasting horizon lengthens, as demonstrated by the widening performance gap between TimesFM and MOIRAI.
+This trend is consistent with our earlier observation that AR-based methods could suffer from significant error accumulation when applied to long-term time-series forecasting.
+Given the inherent strengths of AR decoding, such as its superiority at capturing strong seasonality and its robust performance in certain short-term forecasting scenarios, it is clear that further research is warranted to explore ways to overcome its limitations in long-term forecasting contexts.
+This could potentially unlock new avenues for enhancing the versatility and effectiveness of AR-based time-series foundation models across a broader range of forecasting horizons.
+The Critical Role of Addressing Complex Data Distributions
+Figure 4 illustrates the incremental changes in CRPS among leading probabilistic time-series foundation models, such as MOIRAI and Chronos, compared to the best-performing short-term probabilistic model, CSDI.
+It becomes apparent that in scenarios characterized by complex data distributions, indicated by higher non-Gaussianity scores, the performance decline of MOIRAI in relation to CSDI becomes notably more pronounced.
+This phenomenon may be attributed to MOIRAI’s approach to supporting distributional forecasting, which involves utilizing a mixture of predefined distribution heads.
+While this method is efficient and effective for certain applications, it may lack the expressiveness required to accurately model more complex data distributions.
+Furthermore, these observations underscore that, in specific contexts, foundation models might not be able to fully replace traditional models that have been specifically tailored and trained for particular domains.
+Additionally, the prospect of fine-tuning these foundation models as a remedy is less economically viable, primarily due to their significantly larger size.
+This highlights the importance of not only continuing to refine foundation models to enhance their adaptability and performance across a wide spectrum of data distributions but also recognizing the continued relevance of domain-specific models, especially for handling intricate data distributions where a more nuanced approach may be necessary.
 5 Conclusion
-This paper presents an integrated perspective on data, metrics, and methodology in the field of deep time-series forecasting, facilitated by the ProbTS tool.
-Our findings not only provide valuable insights into the strengths and limitations of existing methods but also highlight several underexplored areas that merit future research.
-In addition, we expect that our insights could spur more explorations on these fields.
-We also hope the ProbTS tool can become a useful resource for the community, expediting future advancements in time-series forecasting.
-Impact Statements
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A More Details on ProbTS
-A.1 Additional Related Work
+In this study, we introduced ProbTS, a benchmark tool tailored for evaluating essential forecasting needs, which facilitates a detailed comparison of various state-of-the-art models in the context of time-series forecasting.
+Through our comprehensive analysis, we identified significant challenges and opportunities in the realm of time-series forecasting, particularly highlighting the need for models that can effectively address both point and probabilistic forecasting across diverse horizons.
+Limitations
+While our study represents a significant step forward in understanding and evaluating time-series forecasting models, it does come with many limitations.
+A predominant focus of our work is on empirical analysis, relying heavily on intuitions and experimental observations, which may lack the depth that theoretical foundations could provide.
+Additionally, our exploration, though extensive, might not encompass all the nuanced factors that influence model performance.
+By concentrating on major methodological decisions such as AR versus NAR decoding schemes, we may inadvertently overlook other critical aspects that could play a decisive role in forecasting accuracy.
+Moreover, the datasets employed for evaluation, despite their diversity and relevance to current research threads, may not fully capture the vast spectrum of real-world forecasting challenges.
+This limitation is particularly pronounced when comparing different foundation models, as their pre-training often involves an even broader array of data, potentially skewing the comparative analysis.
+Future Directions
+The insights derived from our study open the door to numerous promising research directions.
+Addressing the shortcomings of AR and NAR decoding schemes, especially in their application across varying forecasting horizons, emerges as a critical area for future exploration.
+Innovating effective architecture designs that can navigate the intricacies of short-term forecasting challenges and devising efficient methods for long-term probabilistic forecasting stand out as urgent needs.
+For those time-series foundation models, developing mechanisms to combat error accumulation in AR-based methods could be very helpful.
+Equally important is the enhancement of models’ abilities to characterize complex data distributions, which could significantly improve the adaptability and effectiveness of foundation models.
+Beyond these technical endeavors, expanding the scope of datasets used for evaluation to encompass a wider range of real-world scenarios will be crucial for validating the robustness and versatility of future forecasting models.
+Lastly, integrating theoretical insights with empirical findings could provide a more holistic understanding of model behaviors, contributing to the development of more sophisticated and nuanced forecasting solutions.
+Checklist
+  1. 1.
+    For all authors…
+    1. (a)
+      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
+    2. (b)
+      Did you describe the limitations of your work? [Yes] See Sections 5.
+    3. (c)
+      Did you discuss any potential negative societal impacts of your work? [N/A]
+    4. (d)
+      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
+  2. 2.
+    If you are including theoretical results…
+    1. (a)
+      Did you state the full set of assumptions of all theoretical results? [N/A]
+    2. (b)
+      Did you include complete proofs of all theoretical results? [N/A]
+  3. 3.
+    If you ran experiments (e.g. for benchmarks)…
+    1. (a)
+      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)?
+[Yes] We release the ProbTS toolkit, documentation, and running scripts at https://github.com/microsoft/ProbTS.
+The repository includes parameter configurations for benchmarking experiments, ensuring reproducibility of all results presented in the paper.
+    2. (b)
+      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)?
+[Yes] All datasets in this paper are split according to standard practices in previous works.
+We have specified the hyperparameters in Appendix B.3 The detailed configuration file for each baseline is available on our GitHub page.
+    3. (c)
+      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)?
+[Yes] We specify a fixed set of random seeds for each experiment to indicate the error bars.
+We show the mean and standard deviation of evaluation scores collected from multiple runs.
+    4. (d)
+      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix B.3.
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
+    1. (a)
+      If your work uses existing assets, did you cite the creators? [Yes]
+    2. (b)
+      Did you mention the license of the assets? [Yes] See Appendix B.4.
+    3. (c)
+      Did you include any new assets either in the supplemental material or as a URL? [Yes]
+    4. (d)
+      Did you discuss whether and how consent was obtained from people whose data you’re using/curating?
+[Yes] The data utilized in this study is open-source and does not necessitate consent.
+Acquisition details are available on our GitHub page.
+    5. (e)
+      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content?
+[Yes] The data we are using contains no personally identifiable information or offensive content.
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects…
+    1. (a)
+      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
+    2. (b)
+      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
+    3. (c)
+      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
+Appendix A Additional Related Work
+A.1 A Comparison with Traditional Models on Covering Essential Forecasting Needs and methodological decisions
 Table 2 presents a comparative summary of our approach, which adopts an integrated perspective, and representative studies from the existing literature.
@@
 The comparison is based on data scenarios (short-term versus long-term forecasting), primary evaluation metrics (point versus distributional forecasts), and key methodological choices (general or customized neural architecture designs, and autoregressive or non-autoregressive decoding schemes).
-Method Fore. Horizon Metrics Arch. Design Dec. Scheme
-Short Long Point Distr. General Customized Autoreg. Non-autoreg.
-N-BEATS (Oreshkin et al., 2020)
-Autoformer (Wu et al., 2021)
-Informer (Zhou et al., 2021)
-Pyraformer (Liu et al., 2022)
-N-HiTS (Challu et al., 2023)
-LTSF-Linear (Zeng et al., 2023)
-PatchTST (Nie et al., 2023)
-TimesNet (Wu et al., 2023)
-DeepAR (Salinas et al., 2020)
-GP-copula (Salinas et al., 2019)
-LSTM NVP (Rasul et al., 2021b)
-LSTM MAF (Rasul et al., 2021b)
-Trans MAF (Rasul et al., 2021b)
-TimeGrad (Rasul et al., 2021a)
-CSDI (Tashiro et al., 2021)
-SPD (Bilos et al., 2023)
-TSDiff (Kollovieh et al., 2023)
-This Paper
-A.2 Time-series Forecasting Datasets
+Method Pred. Horizon Paradigm Arch. Design Dec. Scheme
+Short Long Point Distr. General Customized AR Non-AR
+N-BEATS [52]
+Autoformer [70]
+Informer [74]
+Pyraformer [38]
+N-HiTS [12]
+LTSF-Linear [72]
+PatchTST [48]
+TimesNet [69]
+iTransformer [39]
+DeepAR [58]
+GP-copula [57]
+LSTM NVP [56]
+LSTM MAF [56]
+Trans MAF [56]
+TimeGrad [55]
+CSDI [60]
+SPD [8]
+TSDiff [33]
+This Study
+A.2 A Comparison of Pre-trained Time-series Foundation Models
+We have incorporated eight recently emerged time series foundation models, namely Lag-Llama [54], Chronos [3], TimesFM [16], Timer [40], MOIRAI [68], UniTS [24], ForecastPFN [19], and TTM [21], into our framework.
+These foundation models are categorized based on their capabilities, such as zero-shot forecasting, adaptability to varying prediction lengths, and support for probabilistic predictions, as well as their architectural designs, including whether they are auto-regressive and the nature of their backbone networks.
+Additionally, we have detailed their training processes, including the lengths of prediction horizons used during pre-training and the sizes of look-back windows.
+These details are summarized in Table 3.
+Furthermore, we have compiled a summary of these foundation models’ pre-training and evaluation on several classical time series forecasting datasets. This compilation is presented in Table 4.
+Table 3: Foundation Models for Time Series.
+Zero-shot indicates whether the original work tests zero-shot capabilities.
+Any-horizon indicates if the same pre-trained model can adapt to prediction tasks of varying lengths.
+AR denotes if the model performs auto-regressive forecasting.
+Prob.
+indicates if the model natively supports probabilistic forecasting.
+Arch.
+denotes the model’s backbone architecture: D-O for decoder-only transformer, E-O for encoder-only transformer, E-D for encoder-decoder transformer, and unique for specially designed backbones.
+Multi-variate indicates if the model explicitly handles multivariate relationships.
+Pre-train Horizons specifies the forecasting task horizons during pre-training.
+Look-back Window specifies the context history length settings used in the original experiments.
+Model Zero-shot Any-horizon AR Prob. Arch. Multi-variate Pre-train Horizons Look-back Window
+Lag-Llama D-O 24$\sim$60 32$\sim$1024
+Chronos E-D 64 512
+TimesFM D-O -- 512
+Timer D-O up to 1440 672
+[…]
+UniTS E-O -- 60$\sim$720
+ForecastPFN E-O 0 50 50 500
+TTM Unique 96 512
+[…]
+Model Solar Wikipedia ETTm1 ETTm2 ETTh1 ETTh2 Electricity Traffic Weather Exchange ILI
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Timer few few few few few few few
+A.3 A Comparison with Existing Libraries on the Coverage of Data, Model, and Metrics
+ProbTS is a research toolkit designed to advance forecasting research across varied horizons, focusing on both point and distributional forecasting.
+To achieve these objectives, ProbTS includes state-of-the-art (SOTA) models, comprehensive evaluation protocols (point vs.
+distributional), and explores different methodological aspects of forecasting models, particularly in terms of distributional estimation methods and decoding schemes (AR vs.
+NAR).
+In Table 5, we provide a comprehensive comparison of ProbTS with existing libraries in terms of toolkit functionalities and the benchmarking aspects we aim to investigate.
+Table 5: Comparison of Various Time Series Tools.
+Tools Features Benchmarking
+SOTA Model Dist. Evaluation Short-term Long-term AR vs. NAR Point vs Prob.
+Darts [28]
+Merlion [7]
+Kats [32]
+pytorch-forecasting [62]
+NeuralForecast [50]
+pytorch-transformer-ts [63]
+Prophet [61]
+sktime [41]
+tsai [51]
+TSlib [69]
+GluonTS [2]
+ProbTS
+Appendix B More Details on ProbTS
+B.1 Data
+The data module unifies varied data scenarios to facilitate thorough evaluation and implements standardized pre-processing techniques to ensure fair comparison.
+Moreover, we utilize a quantitative approach to visually delineate datasets’ intrinsic characteristics, which employs decomposition to assess trends and seasonality in a time series and evaluate the similarity between data distribution and a Gaussian to depict the complexity of data distribution.
+B.1.1 Time-series Forecasting Datasets
 Table 6 provides a summary of the public datasets employed in our study. These datasets have been sourced from recent research studies in the field of deep time-series forecasting.
@@
 Wikipedia 2,000 $\mathbb{N}$ D 792 Page views of 2000 Wikipedia pages
-A.3 Data Visualization
+B.1.2 Data Visualization
 To provide a more tangible understanding of the different forecasting scenarios, we visualize time-series segments from both short-term and long-term forecasting datasets.
@@
 These characteristics render the series more regular patterns in the long-term scenarios.
-It’s important to note that these visualizations are not selectively chosen or ”cherry-picked”.
+It’s important to note that these visualizations are not selectively chosen or "cherry-picked".
 We have depicted multiple time-series segments from various time steps, and the observed patterns remain consistent across these different instances.
@@
 Figure 6: We have also sampled and visualized multiple time-series segments from the long-term forecasting datasets, where the size of the segment window matches the prediction horizon.
-A.4 Quantifying Trend and Seasonality Strengths
-Based on the intuition obtained from data visualization, we would like to quantify the strengths of trend and seasonality for a time-series segment with a predefined window size (corresponding to the prediction horizon length).
+B.1.3 Quantifying Trend and Seasonality Strengths
+Based on the intuition obtained from data visualization, we would like to quantify the strengths of trend and seasonality for a time-series segment with a predefined window size (corresponding to the prediction horizon).
 Then we can quantify the trend and seasonality strengths at the dataset level by averaging over all time-series segments of a dataset.
-To quantify the strengths of trend and seasonality for a fixed-length time-series segment, we draw upon methodologies outlined in the work of (Wang et al., 2006).
+To quantify the strengths of trend and seasonality for a fixed-length time-series segment, we draw upon methodologies outlined in the work of [66].
 In particular, we employed a time series decomposition model expressed as:
@@
 As the forecasting horizon extends, the dataset appears a more pronounced trend while retaining minimal seasonality.
-A.5 Quantifying Data Distribution Complexity
+B.1.4 Quantifying Data Distribution Complexity
 To differentiate between methods optimized for point or distributional forecasts, we aim to quantify the complexity of data distribution within a time-series segment whose window size equals the prediction horizon length.
@@
 In contrast, advanced probabilistic methods, which do not make prior assumptions about data distribution, can adapt to complex data distributions in a data-driven manner.
-Hence, we use the Jensen–Shannon divergence (Nielsen, 2019) to measure the similarity between the actual value distribution of a time-series segment and a Gaussian distribution fitted to the observed values.
+Hence, we use the Jensen–Shannon divergence [49] to measure the similarity between the actual value distribution of a time-series segment and a Gaussian distribution fitted to the observed values.
 Short-term datasets used a window size of 30, while long-term datasets used a size of 336.
@@
 A larger divergence value indicates a larger deviation from a Gaussian distribution in the data.
-A.6 Evaluation Metrics
+B.2 Metrics
 In ProbTS, we integrate an extensive variety of metrics that take into account both point and distributional forecasts, thereby providing a comprehensive and multifaceted assessment of forecasting models.
-A.6.1 Metrics for Point Forecasts
+B.2.1 Metrics for Point Forecasts
 Mean Absolute Error (MAE)
@@
 Its mathematical formula is given by:
-$\textrm{MAE}=\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{T}_{t=1}|x_{i,t}-\hat{x}_{i,t}|,$
-where $K$ is the number of variates, $L$ is the length of series, $x_{i,t}$ and $\hat{x}_{i,t}$ denotes the ground-truth value and the predicted value, respectively.
+$\textrm{MAE}=\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}-\hat{x}^{k}_{t}|,$
+where $K$ is the number of variates, $T$ is the length of series, $x^{k}_{t}$ and $\hat{x}^{k}_{t}$ denotes the ground-truth value and the predicted value, respectively.
 For multivariate time series, we also provide the aggregated version:
 $\textrm{MAE}_{\textrm{sum}}=\frac{1}{T}\sum^{T}_{t=1}|x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t}|,$
-where $x^{\textrm{sum}}_{t}$ and $\hat{x}^{\textrm{sum}}_{t}$ are the summation across the dimension $K$ of $x_{i,t}$ and $\hat{x}_{i,t}$, respectively.
+where $x^{\textrm{sum}}_{t}$ and $\hat{x}^{\textrm{sum}}_{t}$ are the summation across the dimension $K$ of $x^{k}_{t}$ and $\hat{x}^{k}_{t}$, respectively.
 Normalized Mean Absolute Error (NMAE)
@@
 The mathematical representation of NMAE is given by:
-$\textrm{NMAE}=\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{T}_{t=1}\frac{|x_{i,t}-\hat{x}_{i,t}|}{|x_{i,t}|}.$
+$\textrm{NMAE}=\frac{\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}-\hat{x}^{k}_{t}|}{\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}|}.$
 Its aggregated version is:
-$\textrm{NMAE}_{\textrm{sum}}=\frac{1}{T}\sum^{T}_{t=1}\frac{|x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t}|}{|x^{\textrm{sum}}_{t}|}.$
+$\textrm{NMAE}_{\textrm{sum}}=\frac{\sum^{T}_{t=1}|x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t}|}{\sum^{T}_{t=1}|x^{\textrm{sum}}_{t}|}.$
 Mean Squared Error (MSE)
 The Mean Squared Error (MSE) is a quantitative metric used to measure the average squared difference between the observed actual value and forecasts. It is defined mathematically as follows:
-$\textrm{MSE}=\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{L}_{t=1}(x_{i,t}-\hat{x}_{i,t})^{2}.$
+$\textrm{MSE}=\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}(x^{k}_{t}-\hat{x}^{k}_{t})^{2}.$
 For multivariate time series, we also provide the aggregated version:
-$\textrm{MSE}_{\textrm{sum}}=\frac{1}{T}\sum^{L}_{t=1}(x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t})^{2}.$
+$\textrm{MSE}_{\textrm{sum}}=\frac{1}{T}\sum^{T}_{t=1}(x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t})^{2}.$
 Normalized Root Mean Squared Error (NRMSE)
@@
 It can be formally written as:
-$\textrm{NRMSE}=\frac{\sqrt{\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{L}_{t=1}(x_{i,t}-\hat{x}_{i,t})^{2}}}{\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{T}_{t=1}|x_{i,t}|}.$
+$\textrm{NRMSE}=\frac{\sqrt{\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}(x^{k}_{t}-\hat{x}^{k}_{t})^{2}}}{\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}|}.$
 For multivariate time series, we also provide the aggregated version:
-$\textrm{NRMSE}_{\textrm{sum}}=\frac{\sqrt{\frac{1}{T}\sum^{L}_{t=1}(x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t})^{2}}}{\frac{1}{T}\sum^{T}_{t=1}|x^{\textrm{sum}}_{t}|}.$
-A.6.2 Metrics for Distributional Forecasts
+$\textrm{NRMSE}_{\textrm{sum}}=\frac{\sqrt{\frac{1}{T}\sum^{T}_{t=1}(x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t})^{2}}}{\frac{1}{T}\sum^{T}_{t=1}|x^{\textrm{sum}}_{t}|}.$
+Mean Absolute Scaled Error (MASE)
+The Mean Absolute Scaled Error (MASE) divideds the MAE of forecasted values by MAE of the in-sample one-step naive forecast, which is a scale-invariant metrics:
+$\textrm{MASE}=\frac{\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}-\hat{x}^{k}_{t}|}{\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}-x^{k}_{t-1}|}.$
+B.2.2 Metrics for Distributional Forecasts
 Continuous Ranked Probability Score (CRPS)
-The Continuous Ranked Probability Score (CRPS) (Matheson & Winkler, 1976) quantifies the agreement between a cumulative distribution function (CDF) $F$ and an observation $x$, represented as:
+The Continuous Ranked Probability Score (CRPS) [46] quantifies the agreement between a cumulative distribution function (CDF) $F$ and an observation $x$, represented as:
 […]
-where $\mathds{I}{x\leq z}$ denotes the indicator function, equating to one if $x\leq z$ and zero otherwise.
+where $\mathds{I}\{x\leq z\}$ denotes the indicator function, equating to one if $x\leq z$ and zero otherwise.
 Being a proper scoring function, CRPS reaches its minimum when the predictive distribution $F$ coincides with the data distribution.
@@
 $\textrm{CRPS}_{\textrm{sum}}=\mathds{E}_{t}\left[\textrm{CRPS}\left(\hat{F}_{\textrm{sum}}(t),\sum^{K}_{i=1}x^{0}_{i,l}\right)\right].$
-A.7 Implementation Details
-ProbTSwas developed using PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019).
+B.3 Baselines
+To ensure the integrity of the results, ProbTS adheres to a standard implementation process, employing unified data splitting, standardization techniques, and adopting fair settings for hyperparameter tuning across all methods.
+Implementation Details
+ProbTS was developed using PyTorch Lightning [23].
 During training, we sampled 100 batches per epoch and limited training to 50 epochs, using the CRPS metric for checkpointing.
@@
 To enable evaluation of distribution-level metrics, we conducted 100 samplings to calculate metrics on the test set.
-Following the most commonly adopted settings (Zeng et al., 2023; Nie et al., 2023; Wu et al., 2023), in the long-term forecasting context, all of the models are following the same experimental setup with prediction length $T\in\{24,36,48,60\}$ for ILI dataset and $T\in\{96,192,336,720\}$ for other datasets.
+Following the most commonly adopted settings [72, 48, 69], in the long-term forecasting context, all of the models are following the same experimental setup with prediction length $T\in\{24,36,48,60\}$ for ILI-L dataset and $T\in\{96,192,336,720\}$ for other datasets.
 Note that the lookback window here is 96 for all the models, to ensure a fair comparison.
-In the short-term forecasting context, the length of the lookback window is the same as the forecasting horizons, which are 30 for Exchange-S dataset and Wikipedia-S dataset, and 24 for the rest, the same as (Salinas et al., 2019).
+In the short-term forecasting context, the length of the lookback window is the same as the forecasting horizons, which are 30 for Exchange-S dataset and Wikipedia-S dataset, and 24 for the rest, the same as [57].
 Hyper-parameter Tuning
@@
 CSDI learning_rate=0.001, channels=64, emb_time_dim=128, emb_feature_dim=16, num_steps=50, num_heads=8, n_layers=4
-Appendix B Additional Results and Experiments
-B.1 Impact of Data Scale
+Implementation Details on Foundation Models
+We used reference implementations of eight time series foundation models into ProbTS.
+For Lag-Llama [54], we use its official code [^5] https://github.com/time-series-foundation-models/lag-llama and integrate the LagLlamaEstimator with its pre-trained checkpoint.
+The look-back window is uniformly set to 512, irrespective of the forecast horizon.
+The other hyper-parameters are aligned with the recommended settings.
+For Chronos [3], we use its official code [^6] https://github.com/amazon-science/chronos-forecasting and integrate the ChronosPipeline into ProbTS using amazon/chronos-t5 checkpoints (three models were tested: small, base and large).
+Two look-back windows are used during evaluation: 96 and 512.
+We set limit_prediction_length=False to enable it to predict horizons longer than 64.
+However, this may potentially lead to a decrease in predictive performance since the model was only trained to consider prediction lengths of 64 or less during pre-training.
+For TimesFM [16], we modify its official code [^7] https://github.com/google-research/timesfm into ProbTS and load checkpoints from google/timesfm-1.0-200m.
+Look-back window is set to 96 for a fair comparison.
+For Timer [40], we modify its official code [^8] https://github.com/thuml/Large-Time-Series-Model into ProbTS and Timer_67M_UTSD_4G checkpoint downloaded from its repo.
+Look-back window is also set to 96 for a fair comparison.
+For MOIRAI [68], we employ its official code [^9] https://github.com/SalesforceAIResearch/uni2ts and load checkpoints from Salesforce/moirai-1.0-R-base.
+Two look-back windows are utilized during evaluation: 96 and 5000.
+The original experiments suggest that MOIRAI’s forecasting capability can be consistently enhanced by increasing the look-back window.
+Consequently, we have included a 5000 look-back window to test the model’s performance.
+For UniTS [24], we have adapted its official code [^10] https://github.com/abacusai/ForecastPFN into our ProbTS framework and loaded the saved_weights from its repo.
+The look-back window is set to 96 to ensure a fair comparison.
+It is important to note that this checkpoint was originally used for the Zero-Shot New-length Forecasting experiment in the original work (where models are challenged to predict new lengths by adjusting from the trained length, with offsets ranging from 0 to 384), which differs from the objectives of our experiments.
+For ForecastPFN [19], we have integrated its official code [^11] https://github.com/SalesforceAIResearch/uni2ts into our ProbTS framework and have utilized the units_x128_pretrain_checkpoint.
+However, we have encountered some challenges in replicating the performance levels reported in the original paper across various datasets.
+It appears that our experience aligns with the observations made in the Chronos paper [3], specifically as depicted in Figure 5, where ForecastPFN’s performance was not as robust as initially anticipated.
+For TTM [21], we have adapted its official code [^12] https://github.com/ibm-granite/granite-tsfm/blob/main/notebooks/hfdemo/ttm_getting_started.ipynb into our ProbTS framework and have loaded the ibm-granite/granite-timeseries-ttm-v1 checkpoint.
+However, this method does not support arbitrary lengths for forecasting.
+The publicly available model currently supports a forecast length of 96 only, and thus, we have not conducted evaluations at other forecast lengths.
+It is worth noting that, due to time constraints, we did not adjust the context window for all models.
+Instead, we chose 96 as a balanced and fair window size, which might result in suboptimal performance for some models.
+B.4 Data and Code Availability
+We release the ProbTS toolkit, documentation, and running scripts at https://github.com/microsoft/ProbTS under the MIT license.
+The repository includes parameter configurations for benchmarking experiments, ensuring reproducibility of all results presented in the paper.
+Most datasets used in this paper licensed under Creative Commons Attribution 4.0 International (CC BY 4.0), accessible via instructions in the repository.
+Appendix C Overall Comparison Results
+C.1 An Overall Comparison of Traditional Time-series Models on Short-term Forecasting
+Table 9: Results ($\textrm{mean}_{\textrm{std}}$) on short-term forecasting scenarios, each containing five independent runs with different seeds.
+Model Exchange-S Solar-S Electricity-S Traffic-S Wikipedia-S
+CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
+Glob. mean $0.188$ $0.188$ $1.403$ $1.403$ $0.412$ $0.412$ $0.540$ $0.540$ $0.577$ $0.577$
+Batch mean $0.012$ $0.012$ $1.244$ $1.244$ $0.365$ $0.365$ $0.503$ $0.503$ $0.336$ $0.336$
+Linear $0.012_{.001}$ $0.012_{.001}$ $0.704_{.036}$ $0.704_{.036}$ $0.138_{.009}$ $0.138_{.009}$ $0.327_{.032}$ $0.327_{.032}$ $0.874_{.151}$ $0.874_{.151}$
+GRU $0.013_{.002}$ $0.013_{.002}$ $0.594_{.144}$ $0.594_{.144}$ $0.134_{.009}$ $0.134_{.009}$ $0.193_{.002}$ $0.193_{.002}$ $0.394_{.013}$ $0.394_{.013}$
+Transformer $0.016_{.001}$ $0.016_{.001}$ $0.538_{.066}$ $0.538_{.066}$ $0.115_{.005}$ $0.115_{.005}$ $0.204_{.006}$ $0.204_{.006}$ $0.408_{.011}$ $0.408_{.011}$
+N-HiTS $0.012_{.000}$ $0.012_{.000}$ $0.572_{.020}$ $0.572_{.020}$ $0.074_{.003}$ $0.074_{.003}$ $0.193_{.002}$ $0.193_{.002}$ $0.332_{.011}$ $0.332_{.011}$
+NLinear $0.010_{.000}$ $0.010_{.000}$ $0.560_{.002}$ $0.560_{.002}$ $0.083_{.002}$ $0.083_{.002}$ $0.233_{.001}$ $0.233_{.001}$ $0.321_{.001}$ $0.321_{.001}$
+DLinear $0.012_{.001}$ $0.012_{.001}$ $0.547_{.009}$ $0.547_{.009}$ $0.076_{.003}$ $0.076_{.003}$ $0.250_{.002}$ $0.250_{.002}$ $0.412_{.001}$ $0.412_{.001}$
+PatchTST $0.010_{.000}$ $0.010_{.000}$ $0.496_{.002}$ $0.496_{.002}$ $0.067_{.001}$ $0.067_{.001}$ $0.202_{.001}$ $0.202_{.001}$ $0.257_{.001}$ $0.257_{.001}$
+TimesNet $0.011_{.001}$ $0.011_{.001}$ $0.507_{.019}$ $0.507_{.019}$ $0.071_{.002}$ $0.071_{.002}$ $0.205_{.002}$ $0.205_{.002}$ $0.304_{.002}$ $0.304_{.002}$
+iTransformer $0.010_{.000}$ $0.010_{.000}$ $0.496_{.000}$ $0.496_{.000}$ $0.074_{.000}$ $0.074_{.000}$ $0.158_{.000}$ $0.158_{.000}$ $0.262_{.000}$ $0.262_{.000}$
+GRU NVP $0.016_{.003}$ $0.020_{.003}$ $0.396_{.021}$ $0.507_{.022}$ $0.055_{.002}$ $0.073_{.003}$ $0.161_{.006}$ $0.203_{.009}$ $0.282_{.003}$ $0.330_{.003}$
+GRU MAF $0.015_{.001}$ $0.020_{.001}$ $0.386_{.026}$ $0.492_{.027}$ $0.051_{.001}$ $0.067_{.001}$ $0.131_{.006}$ $0.165_{.009}$ $0.281_{.004}$ $0.337_{.005}$
+Trans MAF $0.011_{.001}$ $0.014_{.001}$ $0.400_{.022}$ $0.503_{.022}$ $0.054_{.004}$ $0.071_{.005}$ $0.129_{.004}$ $0.165_{.006}$ $0.289_{.008}$ $0.344_{.008}$
+TimeGrad $0.011_{.001}$ $0.014_{.002}$ $0.359_{.011}$ $0.445_{.023}$ $0.052_{.001}$ $0.067_{.001}$ $0.164_{.091}$ $0.201_{.115}$ $0.272_{.008}$ $0.327_{.011}$
+CSDI $0.008_{.000}$ $0.011_{.000}$ $0.366_{.005}$ $0.484_{.008}$ $0.050_{.001}$ $0.065_{.001}$ $0.146_{.012}$ $0.176_{.013}$ $0.219_{.006}$ $0.259_{.009}$
+Table 9 presents a comprehensive comparison of various time-series models in ProbTS on short-term forecasting scenarios.
+The results, reported as mean ± standard deviation, are derived from five independent runs with different seeds for each scenario.
+C.2 An Overall Comparison of Traditional Time-series Models on Long-term Forecasting
+Table 10 presents a comprehensive comparison of various time-series models in ProbTS on long-term forecasting scenarios.
+The results, reported as mean ± standard deviation, are derived from five independent runs with different seeds for each scenario.
+The input sequence length is set to 36 for the ILI dataset and 96 for the others.
+Due to the excessive time and memory consumption of CSDI in producing long-term forecasts, its results are unavailable in some datasets.
+Table 10: Results ($\textrm{mean}_{\textrm{std}}$) on long-term forecasting scenarios, each containing five independent runs with different seeds.
+The input sequence length is set to 36 for the ILI-L dataset and 96 for the others.
+Due to the excessive time and memory consumption of CSDI in producing long-term forecasts, its results are unavailable in some datasets.
+Dataset Pred iTransformer PatchTST DLinear Autoformer CSDI TimeGrad GRU NVP
+len CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
+ETTm1-L 96 $\underline{0.271_{.000}}$ $\bm{0.271_{.000}}$ $0.272_{.001}$ $\underline{0.272_{.001}}$ $0.282_{.002}$ $0.282_{.002}$ $0.388_{.001}$ $0.388_{.001}$ $\bm{0.236_{.006}}$ $0.308_{.005}$ $0.522_{.105}$ $0.645_{.129}$ $0.383_{.053}$ $0.488_{.058}$
+ETTm1-L 192 $0.301_{.000}$ $\underline{0.301_{.000}}$ $\underline{0.295_{.001}}$ $\bm{0.295_{.001}}$ $0.309_{.004}$ $0.309_{.004}$ $0.442_{.001}$ $0.442_{.001}$ $\bm{0.291_{.025}}$ $0.377_{.026}$ $0.603_{.092}$ $0.748_{.084}$ $0.396_{.030}$ $0.514_{.042}$
+ETTm1-L 336 $0.333_{.000}$ $\underline{0.333_{.000}}$ $\underline{0.323_{.001}}$ $\bm{0.323_{.001}}$ $0.338_{.008}$ $0.338_{.008}$ $0.429_{.000}$ $0.429_{.000}$ $\bm{0.322_{.033}}$ $0.419_{.042}$ $0.601_{.028}$ $0.759_{.015}$ $0.486_{.032}$ $0.630_{.029}$
+ETTm1-L 720 $\underline{0.376_{.000}}$ $\underline{0.376_{.000}}$ $\bm{0.353_{.001}}$ $\bm{0.353_{.001}}$ $0.387_{.006}$ $0.387_{.006}$ $0.440_{.000}$ $0.440_{.000}$ $0.448_{.038}$ $0.578_{.051}$ $0.621_{.037}$ $0.793_{.034}$ $0.546_{.036}$ $0.707_{.050}$
+ETTm2-L 96 $0.137_{.000}$ $\underline{0.137_{.000}}$ $\underline{0.132_{.001}}$ $\bm{0.132_{.001}}$ $0.138_{.000}$ $0.138_{.000}$ $0.158_{.000}$ $0.158_{.000}$ $\bm{0.115_{.009}}$ $0.146_{.012}$ $0.427_{.042}$ $0.525_{.047}$ $0.319_{.044}$ $0.413_{.059}$
+ETTm2-L 192 $0.161_{.000}$ $\underline{0.161_{.000}}$ $\underline{0.157_{.001}}$ $\bm{0.157_{.001}}$ $0.163_{.003}$ $0.163_{.003}$ $0.175_{.000}$ $0.175_{.000}$ $\bm{0.147_{.008}}$ $0.189_{.012}$ $0.424_{.061}$ $0.530_{.060}$ $0.326_{.025}$ $0.427_{.033}$
+ETTm2-L 336 $\underline{0.180_{.000}}$ $\underline{0.180_{.000}}$ $\bm{0.176_{.000}}$ $\bm{0.176_{.000}}$ $0.188_{.001}$ $0.188_{.001}$ $0.191_{.000}$ $0.191_{.000}$ $0.190_{.018}$ $0.248_{.024}$ $0.469_{.049}$ $0.566_{.047}$ $0.449_{.145}$ $0.580_{.169}$
+ETTm2-L 720 $\underline{0.211_{.000}}$ $\underline{0.211_{.000}}$ $\bm{0.205_{.001}}$ $\bm{0.205_{.001}}$ $0.219_{.003}$ $0.219_{.003}$ $0.217_{.000}$ $0.217_{.000}$ $0.239_{.035}$ $0.306_{.040}$ $0.470_{.054}$ $0.561_{.044}$ $0.561_{.273}$ $0.749_{.385}$
+ETTh1-L 96 $\bm{0.321_{.000}}$ $\bm{0.321_{.000}}$ $\underline{0.328_{.003}}$ $\underline{0.328_{.003}}$ $0.352_{.011}$ $0.352_{.011}$ $0.367_{.000}$ $0.367_{.000}$ $0.437_{.018}$ $0.557_{.022}$ $0.455_{.046}$ $0.585_{.058}$ $0.379_{.030}$ $0.481_{.037}$
+ETTh1-L 192 $\bm{0.359_{.000}}$ $\bm{0.359_{.000}}$ $\bm{0.359_{.002}}$ $\bm{0.359_{.002}}$ $0.393_{.001}$ $0.393_{.001}$ $\underline{0.392_{.000}}$ $\underline{0.392_{.000}}$ $0.496_{.051}$ $0.625_{.065}$ $0.516_{.038}$ $0.680_{.058}$ $0.425_{.019}$ $0.531_{.018}$
+ETTh1-L 336 $\underline{0.388_{.000}}$ $\underline{0.388_{.000}}$ $\bm{0.384_{.002}}$ $\bm{0.384_{.002}}$ $0.419_{.007}$ $0.419_{.007}$ $0.398_{.000}$ $0.398_{.000}$ $0.454_{.025}$ $0.574_{.026}$ $0.512_{.026}$ $0.666_{.047}$ $0.458_{.054}$ $0.580_{.064}$
+ETTh1-L 720 $\underline{0.408_{.000}}$ $\underline{0.408_{.000}}$ $\bm{0.397_{.002}}$ $\bm{0.397_{.002}}$ $0.502_{.029}$ $0.502_{.029}$ $0.433_{.000}$ $0.433_{.000}$ $0.528_{.012}$ $0.657_{.014}$ $0.523_{.027}$ $0.672_{.015}$ $0.502_{.039}$ $0.643_{.046}$
+ETTh2-L 96 $\underline{0.177_{.000}}$ $\bm{0.177_{.000}}$ $\underline{0.177_{.000}}$ $\bm{0.177_{.000}}$ $0.211_{.027}$ $0.211_{.027}$ $0.203_{.000}$ $\underline{0.203_{.000}}$ $\bm{0.164_{.013}}$ $0.214_{.018}$ $0.358_{.026}$ $0.448_{.031}$ $0.432_{.141}$ $0.548_{.158}$
+ETTh2-L 192 $\underline{0.203_{.000}}$ $\underline{0.203_{.000}}$ $\bm{0.201_{.001}}$ $\bm{0.201_{.001}}$ $0.238_{.028}$ $0.238_{.028}$ $0.226_{.000}$ $0.226_{.000}$ $0.226_{.018}$ $0.294_{.027}$ $0.457_{.081}$ $0.575_{.089}$ $0.625_{.170}$ $0.766_{.223}$
+ETTh2-L 336 $\underline{0.243_{.000}}$ $\underline{0.243_{.000}}$ $\bm{0.240_{.001}}$ $\bm{0.240_{.001}}$ $0.284_{.008}$ $0.284_{.008}$ $0.264_{.000}$ $0.264_{.000}$ $0.274_{.022}$ $0.353_{.028}$ $0.481_{.078}$ $0.606_{.095}$ $0.793_{.319}$ $0.942_{.408}$
+ETTh2-L 720 $\underline{0.264_{.000}}$ $\underline{0.264_{.000}}$ $\bm{0.252_{.000}}$ $\bm{0.252_{.000}}$ $0.307_{.000}$ $0.307_{.000}$ $0.287_{.000}$ $0.287_{.000}$ $0.302_{.040}$ $0.382_{.030}$ $0.445_{.016}$ $0.550_{.018}$ $0.539_{.090}$ $0.688_{.161}$
+Electricity-L 96 $0.098_{.000}$ $0.098_{.000}$ $\bm{0.086_{.001}}$ $\bm{0.086_{.001}}$ $\underline{0.090_{.001}}$ $\underline{0.090_{.001}}$ $0.140_{.000}$ $0.140_{.000}$ $0.153_{.137}$ $0.203_{.189}$ $0.096_{.002}$ $0.119_{.003}$ $0.094_{.003}$ $0.118_{.003}$
+Electricity-L 192 $0.106_{.000}$ $0.106_{.000}$ $\bm{0.092_{.001}}$ $\bm{0.092_{.001}}$ $\underline{0.095_{.001}}$ $\underline{0.095_{.001}}$ $0.136_{.000}$ $0.136_{.000}$ $0.200_{.094}$ $0.264_{.129}$ $0.100_{.004}$ $0.124_{.005}$ $0.097_{.002}$ $0.121_{.003}$
+Electricity-L 336 $0.115_{.000}$ $0.115_{.000}$ $\underline{0.100_{.000}}$ $\bm{0.100_{.000}}$ $0.104_{.000}$ $\underline{0.104_{.000}}$ $0.147_{.000}$ $0.147_{.000}$- -- $0.102_{.007}$ $0.126_{.008}$ $\bm{0.099_{.001}}$ $0.123_{.001}$
+Electricity-L 720 $0.133_{.000}$ $0.133_{.000}$ $0.116_{.000}$ $\bm{0.116_{.000}}$ $0.122_{.001}$ $\underline{0.122_{.001}}$ $0.159_{.000}$ $0.159_{.000}$- -- $\bm{0.108_{.003}}$ $0.134_{.004}$ $\underline{0.114_{.013}}$ $0.144_{.017}$
+Traffic-L 96 $0.246_{.000}$ $0.246_{.000}$ $0.248_{.001}$ $0.248_{.001}$ $0.356_{.009}$ $0.356_{.009}$ $0.293_{.000}$ $0.293_{.000}$- -- $\underline{0.202_{.004}}$ $\underline{0.234_{.006}}$ $\bm{0.187_{.002}}$ $\bm{0.231_{.003}}$
+Traffic-L 192 $0.259_{.000}$ $0.259_{.000}$ $0.245_{.001}$ $0.245_{.001}$ $0.346_{.009}$ $0.346_{.009}$ $0.318_{.000}$ $0.318_{.000}$- -- $\underline{0.208_{.003}}$ $\underline{0.239_{.004}}$ $\bm{0.192_{.001}}$ $\bm{0.236_{.002}}$
+Traffic-L 336 $0.283_{.000}$ $0.283_{.000}$ $0.257_{.002}$ $0.257_{.002}$ $0.350_{.008}$ $0.350_{.008}$ $0.332_{.000}$ $0.332_{.000}$- -- $\underline{0.213_{.003}}$ $\bm{0.246_{.003}}$ $\bm{0.201_{.004}}$ $\underline{0.248_{.006}}$
+Traffic-L 720 $0.275_{.000}$ $0.275_{.000}$ $0.266_{.001}$ $0.266_{.001}$ $0.365_{.009}$ $0.365_{.009}$ $0.341_{.003}$ $0.341_{.003}$- -- $\underline{0.220_{.002}}$ $\bm{0.263_{.001}}$ $\bm{0.211_{.004}}$ $\underline{0.264_{.006}}$
+Weather-L 96 $0.089_{.000}$ $\underline{0.089_{.000}}$ $\underline{0.087_{.002}}$ $\bm{0.087_{.002}}$ $0.112_{.001}$ $0.112_{.001}$ $0.239_{.004}$ $0.239_{.004}$ $\bm{0.068_{.008}}$ $\bm{0.087_{.012}}$ $0.130_{.017}$ $0.164_{.023}$ $0.116_{.013}$ $0.145_{.017}$
+Weather-L 192 $0.093_{.000}$ $0.093_{.000}$ $\underline{0.090_{.001}}$ $\underline{0.090_{.001}}$ $0.122_{.001}$ $0.122_{.001}$ $0.213_{.000}$ $0.213_{.000}$ $\bm{0.068_{.006}}$ $\bm{0.086_{.007}}$ $0.127_{.019}$ $0.158_{.024}$ $0.122_{.021}$ $0.147_{.025}$
+Weather-L 336 $0.096_{.000}$ $\underline{0.096_{.000}}$ $\underline{0.092_{.002}}$ $\bm{0.092_{.002}}$ $0.130_{.002}$ $0.130_{.002}$ $0.176_{.000}$ $0.176_{.000}$ $\bm{0.083_{.002}}$ $0.098_{.002}$ $0.130_{.006}$ $0.162_{.006}$ $0.128_{.011}$ $0.160_{.012}$
+Weather-L 720 $0.099_{.000}$ $\underline{0.099_{.000}}$ $\underline{0.094_{.001}}$ $\bm{0.094_{.001}}$ $0.144_{.001}$ $0.144_{.001}$ $0.170_{.001}$ $0.170_{.001}$ $\bm{0.087_{.003}}$ $0.102_{.005}$ $0.113_{.011}$ $0.136_{.020}$ $0.110_{.004}$ $0.135_{.008}$
+Exchange-L 96 $0.025_{.000}$ $0.025_{.000}$ $\bm{0.023_{.000}}$ $\bm{0.023_{.000}}$ $\underline{0.024_{.000}}$ $\underline{0.024_{.000}}$ $0.032_{.000}$ $0.032_{.000}$ $0.028_{.003}$ $0.036_{.005}$ $0.068_{.003}$ $0.079_{.002}$ $0.071_{.006}$ $0.091_{.009}$
+Exchange-L 192 $0.036_{.000}$ $0.036_{.000}$ $\bm{0.034_{.000}}$ $\bm{0.034_{.000}}$ $\underline{0.035_{.000}}$ $\underline{0.035_{.000}}$ $0.041_{.000}$ $0.041_{.000}$ $0.045_{.003}$ $0.058_{.005}$ $0.087_{.013}$ $0.100_{.019}$ $0.068_{.004}$ $0.087_{.005}$
+Exchange-L 336 $\bm{0.048_{.000}}$ $\bm{0.048_{.000}}$ $\bm{0.048_{.000}}$ $\bm{0.048_{.000}}$ $\bm{0.048_{.001}}$ $\bm{0.048_{.001}}$ $\underline{0.056_{.000}}$ $\underline{0.056_{.000}}$ $0.060_{.004}$ $0.076_{.006}$ $0.074_{.009}$ $0.086_{.008}$ $0.072_{.002}$ $0.091_{.002}$
+Exchange-L 720 $0.076_{.000}$ $0.076_{.000}$ $\bm{0.072_{.000}}$ $\bm{0.072_{.000}}$ $\underline{0.075_{.002}}$ $\underline{0.075_{.002}}$ $0.112_{.002}$ $0.112_{.002}$ $0.143_{.020}$ $0.173_{.020}$ $0.099_{.015}$ $0.113_{.016}$ $0.079_{.009}$ $0.103_{.009}$
+ILI-L 24 $\bm{0.094_{.000}}$ $\bm{0.094_{.000}}$ $0.169_{.005}$ $0.169_{.005}$ $0.213_{.038}$ $0.213_{.038}$ $\underline{0.122_{.000}}$ $\underline{0.122_{.000}}$ $0.250_{.013}$ $0.263_{.012}$ $0.275_{.047}$ $0.296_{.044}$ $0.257_{.003}$ $0.283_{.001}$
+ILI-L 36 $\bm{0.102_{.000}}$ $\bm{0.102_{.000}}$ $0.156_{.005}$ $0.156_{.005}$ $0.230_{.015}$ $0.230_{.015}$ $\underline{0.111_{.000}}$ $\underline{0.111_{.000}}$ $0.285_{.010}$ $0.298_{.011}$ $0.272_{.057}$ $0.298_{.048}$ $0.281_{.004}$ $0.307_{.007}$
+ILI-L 48 $\bm{0.103_{.000}}$ $\bm{0.103_{.000}}$ $0.156_{.008}$ $0.156_{.008}$ $0.221_{.009}$ $0.221_{.009}$ $\underline{0.134_{.000}}$ $\underline{0.134_{.000}}$ $0.285_{.036}$ $0.301_{.034}$ $0.295_{.033}$ $0.320_{.025}$ $0.288_{.008}$ $0.314_{.009}$
+ILI-L 60 $\bm{0.128_{.000}}$ $\bm{0.128_{.000}}$ $0.147_{.003}$ $0.147_{.003}$ $0.230_{.013}$ $0.230_{.013}$ $\underline{0.144_{.000}}$ $\underline{0.144_{.000}}$ $0.283_{.012}$ $0.299_{.013}$ $0.295_{.083}$ $0.325_{.068}$ $0.307_{.005}$ $0.333_{.005}$
+C.3 An Overall Comparison of Time-series Foundation Models on Diverse Prediction Horizons
+Based on the results in Table 4, we selected several datasets that most foundation models have not been pre-trained on for zero-shot evaluation.
+We compared these foundation models with fully-supervised traditional non-universal time-series models.
+The results, presented in Table 11, are reported as normalized MAE (NMAE).
+To comprehensively evaluate short-term and long-term forecasting performance, we selected prediction horizons of $\{24,48,96,192,336,720\}$.
+Due to time constraints, the context window for most models was set to 96 unless otherwise specified.
+For Lag-Llama, we chose a context window of 512 based on explicit recommendations from the original paper and model specifications.
+Additionally, on the MOIRAI model, we explored the impact of longer context windows.
+In some datasets (e.g., ETTh2-L, Weather-L), there were significant improvements, which we retained for reference.
+Other models also utilized longer context windows, but without consistent performance gains.
+It is worth noting that the longer context window of MOIRAI had an adverse effect on the Electricity dataset.
+To achieve optimal performance on unseen data, these models may require a hyperparameter search using validation data, which we leave for future work.
+Table 11: Results of time-series foundation models on diverse prediction horizons.
+Mean NMAE value of five independent runs with different seeds is reported.
+The input sequence length is set to 96 if not specified.
+For every model, we exclude the evaluation results on its pre-trained datasets
+Dataset Pred MOIRAI-5000 MOIRAI Lag-Llama-512 Chronos TimesFM Timer UniTS ForecastPFN CSDI DLinear PatchTST iTransformer
+ETTm1-L 24 0.144 0.128 0.231 0.113 0.133 0.182 0.353 0.878 0.115 0.121 0.199 0.116
+ETTm1-L 48 0.269 0.343 0.244 0.307 0.269 0.339 0.432 0.997 0.266 0.238 0.199 0.243
+ETTm1-L 96 0.296 0.449 0.399 0.393 0.324 0.384 0.457 0.961 0.303 0.283 0.271 0.269
+ETTm1-L 192 0.311 0.479 0.416 0.422 0.378 0.423 0.466 1.031 0.389 0.309 0.295 0.304
+ETTm1-L 336 0.321 0.471 0.429 0.439 0.435 0.446 0.475 1.012 0.449 0.337 0.324 0.337
+ETTm1-L 720 0.351 0.515 0.472 0.467 0.511 0.480 0.499 1.053 0.530 0.385 0.353 0.380
+ETTm2-L 24 0.100 0.113 0.128 0.110 0.117 0.130 0.173 1.506 0.096 0.103 0.162 0.106
+ETTm2-L 48 0.116 0.132 0.162 0.132 0.132 0.141 0.165 1.471 0.121 0.121 0.162 0.118
+ETTm2-L 96 0.141 0.168 0.188 0.158 0.156 0.153 0.169 1.386 0.133 0.138 0.133 0.141
+ETTm2-L 192 0.159 0.192 0.205 0.183 0.186 0.175 0.184 1.397 0.201 0.167 0.158 0.158
+ETTm2-L 336 0.173 0.209 0.226 0.206 0.209 0.193 0.199 1.422 0.226 0.188 0.176 0.184
+ETTm2-L 720 0.200 0.242 0.249 0.240 0.246 0.220 0.223 1.485 0.264 0.222 0.206 0.213
+ETTh1-L 24 0.304 0.297 0.313 0.265 0.277 0.315 0.453 1.141 0.292 0.277 0.356 0.275
+ETTh1-L 48 0.312 0.326 0.341 0.294 0.307 0.339 0.461 1.161 0.392 0.311 0.356 0.304
+ETTh1-L 96 0.324 0.354 0.353 0.321 0.332 0.361 0.469 1.157 0.534 0.340 0.326 0.319
+ETTh1-L 192 0.350 0.393 0.376 0.375 0.383 0.399 0.482 1.226 0.698 0.394 0.357 0.364
+ETTh1-L 336 0.366 0.417 0.393 0.410 0.411 0.433 0.500 1.183 0.603 0.423 0.383 0.389
+ETTh1-L 720 0.380 0.448 0.424 0.432 0.418 0.460 0.499 1.037 0.664 0.501 0.399 0.402
+ETTh2-L 24 0.126 0.142 0.160 0.127 0.134 0.146 0.197 1.650 0.133 0.146 0.213 0.135
+ETTh2-L 48 0.147 0.170 0.195 0.161 0.166 0.166 0.199 1.667 0.186 0.176 0.213 0.161
+ETTh2-L 96 0.162 0.200 0.203 0.194 0.190 0.179 0.204 1.685 0.204 0.209 0.176 0.177
+ETTh2-L 192 0.189 0.244 0.220 0.225 0.220 0.205 0.223 1.768 0.272 0.206 0.200 0.205
+ETTh2-L 336 0.224 0.269 0.245 0.252 0.266 0.247 0.262 1.719 0.321 0.293 0.240 0.244
+ETTh2-L 720 0.243 0.293 0.248 0.290 0.277 0.254 0.264 1.542 0.417 0.307 0.251 0.263
+Weather-L 24 0.043 0.060 0.062 0.060 -- 0.063 -- 1.916 0.050 0.095 0.083 0.074
+Weather-L 48 0.066 0.110 0.086 0.128 -- 0.098 -- 1.924 0.093 0.125 0.156 0.090
+Weather-L 96 0.074 0.134 0.096 0.163 -- 0.109 -- 1.924 0.092 0.113 0.086 0.089
+Weather-L 192 0.074 0.126 0.103 0.161 -- 0.116 -- 1.926 0.079 0.121 0.089 0.099
+Weather-L 336 0.075 0.129 0.109 0.177 -- 0.121 -- 1.930 0.100 0.131 0.092 0.100
+Weather-L 720 0.076 0.151 0.120 0.208 -- 0.124 -- 1.970 0.108 0.145 0.094 0.111
+Electricity-L 24 0.227 0.091 -- -- -- 0.114 -- -- -- 0.093 -- 0.085
+Electricity-L 48 0.210 0.100 -- -- -- 0.128 -- -- -- 0.097 -- 0.089
+Electricity-L 96 0.194 0.101 -- -- -- 0.130 -- -- 0.153 0.090 0.086 0.098
+Electricity-L 192 0.200 0.107 -- -- -- 0.147 -- -- 0.200 0.095 0.092 0.106
+Electricity-L 336 0.202 0.119 -- -- -- 0.168 -- -- -- 0.104 0.100 0.115
+Electricity-L 720 0.217 0.190 -- -- -- 0.205 -- -- -- 0.121 0.116 0.133
+Table 12: Results of probabilistic foundation models on short-term distributional forecasting. For every model, we exclude the evaluation results on its pre-trained datasets.
+Model Exchange-S Solar-S Electricity-S Traffic-S
+CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
+CSDI $0.008_{.000}$ $0.011_{.000}$ $0.366_{.005}$ $0.484_{.008}$ $0.050_{.001}$ $0.065_{.001}$ $0.146_{.012}$ $0.176_{.013}$
+Chronos $0.007$ $0.010$- -- -- -- $0.178$ $0.211$
+MOIRAI-96 $0.007$ $0.010$ $0.502$ $0.681$ $0.069$ $0.084$- --
+MOIRAI-5000 $0.007$ $0.010$ $0.521$ $0.702$ $0.059$ $0.079$- --
+C.4 An Overall Comparison of Time-series Foundation Models on Short-term Probabilistic Forecasting
+Table 12 presents a comparison of two time-series probabilistic foundation models in short-term forecasting scenarios.
+We excluded the results of datasets that has been used in pre-training based on table 4.
+We also explored both short and long context windows for MOIRAI.
+Appendix D Additional Results and Experiments
+D.1 Impact of Data Scale
 To further explore critical characteristics of time-series forecasting, we have examined the correlation between model performance gains, relative to the baseline model (GRU), and dataset dimensions, length, and volume (see Table 13).
@@
 # Var. $\times$ Timestep 0.0162 0.0162 0.0166 0.0166 -0.0068 -0.0011 0.0082 0.0117 -0.0053 -0.0133
-B.2 Statistical and Gradient Boosting Decision Tree Baselines
-To enhance the empirical robustness of our study, we integrate classical statistical models, including ARIMA (Makridakis & Hibon, 1997) and ETS (Hyndman & Athanasopoulos, 2018), along with the Gradient Boosting Decision Tree (GBDT) model, XGBoost, into the ProbTSframework.
+D.2 Statistical and Gradient Boosting Decision Tree Baselines
+To enhance the empirical robustness of our study, we integrate classical statistical models, including ARIMA [43] and ETS [31], along with the Gradient Boosting Decision Tree (GBDT) model, XGBoost, into the ProbTS framework.
 The results in Table 14 clearly demonstrate the superior performance of deep learning methods over simple statistical baselines, emphasizing the importance of capturing non-linear dependencies for accurate forecasts.
@@
 Conversely, in cases of strong trending and weak seasonality, as observed in the ’Wikipedia’ dataset, ARIMA significantly outperforms ETS.
-Utilizing the implementation from (Elsayed et al., 2021), we find that XGBoost competes well, even surpassing neural network models in certain scenarios.
+Utilizing the implementation from [22], we find that XGBoost competes well, even surpassing neural network models in certain scenarios.
 However, for datasets with more complex distributions like ’Solar’ and ’Electricity,’ advanced probabilistic estimation methods demonstrate a substantial advantage over traditional learning methods and point estimation techniques.
@@
 Table 14: Results of statistical models and GBDT baseline on short-term forecasting datasets.
-Model Exchange Rate Solar Electricity Traffic Wikipedia
+Model Exchange-S Solar-S Electricity-S Traffic-S Wikipedia-S
 CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
@@
 CSDI $0.008_{.000}$ $0.011_{.000}$ $0.366_{.005}$ $0.484_{.008}$ $0.050_{.001}$ $0.065_{.001}$ $0.146_{.012}$ $0.176_{.013}$ $0.219_{.006}$ $0.259_{.009}$
-B.3 Experiments on Univariate Datasets
-In pursuit of a comprehensive analysis spanning univariate and multivariate scenarios, we examined a subset of M4 (Makridakis et al., 2020), M5 (Makridakis et al., 2022), and TOURISM datasets (Athanasopoulos et al., 2011)—crucial datasets for univariate time-series forecasting.
+D.3 Experiments on Univariate Datasets
+In pursuit of a comprehensive analysis spanning univariate and multivariate scenarios, we examined a subset of M4 [44], M5 [45], and TOURISM datasets [4]—crucial datasets for univariate time-series forecasting.
 Table 15 provides a quantitative assessment of the intrinsic characteristics of these new datasets, focusing on trending strength, seasonality, and data distribution complexity, as detailed in our paper.
@@
 Probabilistic estimation methods like GRU NVP and TimeGrad excel on datasets with complex distributions (e.g., M4-Weekly and M5), while simpler point forecasting methods such as DLinear and PatchTST perform well on datasets with relatively simple data distribution, like TOURISM-Monthly.
-Both autoregressive and non-autoregressive decoding schemes show comparable performance in short-term forecasting, as discussed in the main paper.”
+Both autoregressive and non-autoregressive decoding schemes show comparable performance in short-term forecasting, as discussed in the main paper."
 Table 16: Results on M4, M5, and TOURISM datasets. We utilize a lookback window of 3H, with ’H’ denoting the forecasting horizon.
-Model DLinear PatchTST GRU NVP TimeGrad
+Dataset DLinear PatchTST GRU NVP TimeGrad
 CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
@@
 TOURISM-Monthly 0.168 0.168 0.136 0.136 0.171 0.223 0.152 0.191
-B.4 Experiments on Synthetic Datasets
+D.4 Experiments on Synthetic Datasets
 To enhance the rigor of the insights presented, we employ synthetic datasets created with the GluonTS library [^13] https://ts.gluon.ai/stable/tutorials/data_manipulation/index.html, encompassing a baseline dataset and variants with pronounced trends, strong seasonality, and complex data distribution (see Table 17).
@@
 CSDI $0.013$ $0.014$ $0.010$ $0.007$ $0.020$ $0.027$ $0.269$ $0.301$
-B.5 Case Study
+D.5 Case Study
 To intuitively demonstrate the distinct characteristics of point and probabilistic estimations, a case study was conducted on short-term datasets.
@@
 Figure 8: Forecasting intervals from the TimeGrad model and the ground-truth value on short-term forecasting datasets.
-B.6 Model Efficiency
+D.6 Model Efficiency
 For reference, detailed results regarding memory usage and time efficiency for five representative models on long-term forecasting datasets are provided here.
@@
 NPARAMS (MB) ETTm1 0.075 2.145 1.079 1.233 1.720
-Electricity 0.076 2.146 3.680 3.472 1.370
-Traffic 0.078 2.149 15.926 8.298 1.390
-Weather 0.075 2.145 3.085 0.574 1.721
-Exchange 0.075 0.135 1.979 0.488 1.720
+Electricity-L 0.076 2.146 3.680 3.472 1.370
+Traffic-L 0.078 2.149 15.926 8.298 1.390
+Weather-L 0.075 2.145 3.085 0.574 1.721
+Exchange-L 0.075 0.135 1.979 0.488 1.720
 Max GPU Mem. (GB) ETTm1 0.002 0.009 0.010 0.012 0.027
-Electricity 0.060 0.068 0.129 0.128 1.411
-Traffic 0.161 0.168 0.361 0.333 9.102
-Weather 0.004 0.012 0.021 0.012 0.070
-Exchange 0.002 0.002 0.013 0.008 0.030
+Electricity-L 0.060 0.068 0.129 0.128 1.411
+Traffic-L 0.161 0.168 0.361 0.333 9.102
+Weather-L 0.004 0.012 0.021 0.012 0.070
+Exchange-L 0.002 0.002 0.013 0.008 0.030
 Table 20: Comparison of inference time (sec./sample).
-Model pred len DLinear PatchTST LSTM NVP TimeGrad CSDI
-ETTm1 96 0.0003 ± 0.0000 0.0003 ± 0.0000 0.0352 ± 0.0007 4.1067 ± 0.0504 16.3280 ± 0.0747
+Dataset Pred len DLinear PatchTST LSTM NVP TimeGrad CSDI
+ETTm1-L 96 0.0003 ± 0.0000 0.0003 ± 0.0000 0.0352 ± 0.0007 4.1067 ± 0.0504 16.3280 ± 0.0747
 192 0.0003 ± 0.0000 0.0003 ± 0.0000 0.0697 ± 0.0020 7.8979 ± 0.0403 25.8378 ± 0.3124
@@
 720 0.0004 ± 0.0000 0.0003 ± 0.0000 0.2603 ± 0.0020 28.6074 ± 1.1346 86.1862 ± 0.1863
-Electricity 96 0.0004 ± 0.0000 0.0045 ± 0.0001 0.1783 ± 0.0006 13.8439 ± 0.0054 388.3150 ± 0.2155
+Electricity-L 96 0.0004 ± 0.0000 0.0045 ± 0.0001 0.1783 ± 0.0006 13.8439 ± 0.0054 388.3150 ± 0.2155
 192 0.0006 ± 0.0000 0.0046 ± 0.0000 0.3700 ± 0.0010 27.6683 ± 0.0368 659.4284 ± 0.2003
@@
 720 0.0015 ± 0.0000 0.0057 ± 0.0000 2.0785 ± 0.0186 104.1473 ± 0.1465 -
-Traffic 96 0.0010 ± 0.0001 0.0102 ± 0.0000 0.3695 ± 0.0022 31.7644 ± 0.0101 -
+Traffic-L 96 0.0010 ± 0.0001 0.0102 ± 0.0000 0.3695 ± 0.0022 31.7644 ± 0.0101 -
 192 0.0013 ± 0.0000 0.0106 ± 0.0000 0.8287 ± 0.0094 63.5832 ± 0.0060 -
@@
 720 0.0039 ± 0.0000 0.0137 ± 0.0000 5.0963 ± 0.0018 258.1274 ± 0.6088 -
-Weather 96 0.0002 ± 0.0000 0.0004 ± 0.0000 $0$ 0.0016 4.1261 ± 0.0812 37.8984 ± 0.0782
+Weather-L 96 0.0002 ± 0.0000 0.0004 ± 0.0000 $0$ 0.0016 4.1261 ± 0.0812 37.8984 ± 0.0782
 192 0.0003 ± 0.0000 0.0004 ± 0.0000 0.1568 ± 0.0008 8.2913 ± 0.5544 62.0223 ± 0.2329
@@
 720 0.0003 ± 0.0000 0.0005 ± 0.0000 0.5447 ± 0.0249 29.4407 ± 0.3519 216.6044 ± 0.4253
-Exchange 96 0.0006 ± 0.0000 0.0004 ± 0.0000 0.0284 ± 0.0001 4.1069 ± 0.0981 17.8655 ± 0.1282
+Exchange-L 96 0.0006 ± 0.0000 0.0004 ± 0.0000 0.0284 ± 0.0001 4.1069 ± 0.0981 17.8655 ± 0.1282
 192 0.0007 ± 0.0000 0.0004 ± 0.0000 0.0563 ± 0.0008 8.1576 ± 0.0911 28.5456 ± 0.0873
@@
 720 0.0007 ± 0.0000 0.0004 ± 0.0000 0.2085 ± 0.0046 30.1443 ± 0.5378 97.7417 ± 0.2606
-ILI 24 0.0002 ± 0.0000 0.0008 ± 0.0001 0.0080 ± 0.0001 1.0427 ± 0.0190 12.4038 ± 0.1681
+ILI-L 24 0.0002 ± 0.0000 0.0008 ± 0.0001 0.0080 ± 0.0001 1.0427 ± 0.0190 12.4038 ± 0.1681
 192 0.0002 ± 0.0000 0.0008 ± 0.0000 0.0121 ± 0.0003 1.5762 ± 0.0282 12.7187 ± 0.1344
@@
 720 0.0002 ± 0.0000 0.0008 ± 0.0000 0.0196 ± 0.0004 2.5787 ± 0.0594 12.5407 ± 0.0481
-B.7 Further Discussion on Cross-channel Interactions
+D.7 Further Discussion on Cross-channel Interactions
 We compile a summary table (Table 21) delineating how models from each branch address the multivariate aspect.
@@
 Model Research branch Process channels independently
-Customized neural architectures N-BEATS (Oreshkin et al., 2020)
-N-HiTS (Challu et al., 2023)
-Autoformer (Wu et al., 2021)
-Informer (Zhou et al., 2021)
-LTSF-Linear (Zeng et al., 2023) ✗/ ✓
-PatchTST (Nie et al., 2023) ✗/ ✓
-TimesNet (Wu et al., 2023)
-Probabilistic estimation DeepAR (Salinas et al., 2020)
-GP-copula (Salinas et al., 2019)
-LSTM NVP (Rasul et al., 2021b)
-LSTM MAF (Rasul et al., 2021b)
-Trans MAF (Rasul et al., 2021b)
-TimeGrad (Rasul et al., 2021a)
-CSDI (Tashiro et al., 2021)
-SPD (Bilos et al., 2023)
+Customized neural architectures N-BEATS [52]
+N-HiTS [12]
+Autoformer [70]
+Informer [74]
+LTSF-Linear [72] ✗/ ✓
+PatchTST [48] ✗/ ✓
+TimesNet [69]
+Probabilistic estimation DeepAR [58]
+GP-copula [57]
+LSTM NVP [56]
+LSTM MAF [56]
+Trans MAF [56]
+TimeGrad [55]
+CSDI [60]
+SPD [8]